### v1

# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

Abstract
Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance. However, achieving differential privacy (DP) in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. In this work, we present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, both with and without secure summation over each node’s direct neighbors. Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage across these scenarios. This framework significantly improves upon prior analyses, for example, reducing the Rényi DP parameter growth from $O(T^{2})$ to $O(T)$, where $T$ is the number of training rounds. We validate our analysis with numerical results demonstrating superior DP bounds compared to existing approaches. We further illustrate our analysis with a logistic regression experiment on MNIST image classification in a fully decentralized setting, demonstrating utility comparable to central aggregation methods.
1 Introduction
Common federated learning (FL) scenarios assume the presence of a central parameter server for coordinating model updates. In contrast, fully decentralized setups operate without a central orchestrator: compute nodes, each holding a private dataset, directly exchange model states or updates with a subset of peers. Such decentralized architectures offer advantages in scalability, fault tolerance, and robustness, but also introduce new algorithmic and privacy challenges.
Fully decentralized gradient-based optimization methods are typically distributed variants of gradient descent and can be broadly classified into two categories.
In random walk-based methods (Lopes and Sayed,, 2007; Johansson et al.,, 2009; Mao et al.,, 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
In gossip averaging-based methods (Boyd et al.,, 2006; Dimakis et al.,, 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
DP in Centralized FL. In centralized FL (Kairouz et al., 2021b, ), distributed DP (Ullah et al.,, 2023; Hartmann and Kairouz,, 2023; Kairouz et al., 2021a, ) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al.,, 2019; Erlingsson et al.,, 2019; Bell et al.,, 2020). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al.,, 2025; Zhu et al.,, 2019; Mrini et al.,, 2024; Pasquini et al.,, 2023). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al.,, 2015; Bellet et al.,, 2018; Xu et al.,, 2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al.,, 2012).
To address this, recent works propose pairwise network DP (Cyffers et al.,, 2024, 2022; Cyffers and Bellet,, 2022), where privacy loss is analyzed between pairs of nodes. These relaxations yield better utility while accounting for the structure of decentralized communication. However, for the practical case of gossip averaging, there does not yet exist a satisfactory composition analysis for the pairwise network DP guarantees. To the best of our knowledge, the best bounds in this case are the Rényi differential privacy (RDP) bounds by Cyffers et al., (2022) which exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of training iterations, making them unsuitable for practical ML model training with reasonable privacy guarantees.
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows empirically as $O(\sqrt{T})$, resulting in $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to considering the threat model of (Cyffers et al.,, 2022), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
Secure Aggregation in Decentralized Learning. While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (Cyffers et al.,, 2022), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can yield a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al.,, 2018; Lian et al.,, 2018; Jeon et al.,, 2021). Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al.,, 2022; Pereira et al.,, 2024; Biswas et al.,, 2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al.,, 2024).
Our Contributions:
  • We develop a novel analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms. Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
  • We demonstrate that the DP guarantees, for both non-adaptive and adaptive compositions-with or without secure summation-can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
  • Our empirical results show that the squared sensitivity scales as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on Rényi differential privacy (RDP).
  • We validate our findings on a decentralized logistic regression task using the MNIST dataset, showing privacy-utility trade-offs comparable to those of central FL.
2 Background
We first shortly review the required technicalities on differential privacy. We then discuss the projected Gaussian mechanism central to our analysis and define the model for decentralized learning and the network DP guarantees.
2.1 Differential Privacy
An input dataset containing $n$ data points is denoted as $D=(x_{1},\ldots,x_{n})\in\mathcal{D}$, where $\mathcal{D}$ denotes the set of datasets of all sizes. We say that two datasets $D$ and $D^{\prime}$ are neighbors if we get one by adding or removing one element to/from the other (denoted $D\sim D^{\prime}$). We say that a mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable.
Definition 1 (Dwork et al., 2006).
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{P}(\mathcal{O})$, where $\mathcal{P}(\mathcal{O})$ denotes the set of probability distributions over the output space $\mathcal{O}$, is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D,D^{\prime}$, every measurable set $O\subset\mathcal{O}$,
$\mathbb{P}(\mathcal{M}(D)\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(\mathcal{M}(D^{\prime})\in O)+\delta.$
Given two distributions $(P,Q)$, if the conditions $\mathbb{P}(P\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(Q\in O)+\delta$ and $\mathbb{P}(Q\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(P\in O)+\delta$ hold for every measurable set $O\subset\mathcal{O}$, we also denote $P\simeq_{(\varepsilon,\delta)}Q$.
With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions. For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $[z]_{+}=\max\{0,z\}$. The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
Lemma 2 (Balle et al., 2018).
A mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if and only if, $\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\delta$.
The Gaussian Mechanism is a common way to achieve $(\epsilon,\delta)$-differential privacy by adding Gaussian noise to a functions output.
Definition 3 (Gaussian Mechanism).
Let $f:\mathcal{D}\to\mathbb{R}^{d}$ be a function with $\ell_{2}$-sensitivity defined as $\Delta_{2}(f)=\max_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$, where the maximum is over all adjacent datasets $D,D^{\prime}$ (i.e., datasets differing in at most one individual’s data). The Gaussian mechanism outputs
$\mathcal{M}(D)=f(D)+Z,$
where $Z\sim\mathcal{N}(0,\sigma^{2}I_{d})$.
Running a DP training algorithm for $T$ iterations is commonly modeled as an adaptive composition of $T$ mechanisms such that the adversary has a view on the output of all intermediate outputs. This means that we then analyze mechanisms of the form
$\mathcal{M}^{(T)}(D)=\big{(}\mathcal{M}_{1}(D),\mathcal{M}_{2}(\mathcal{M}_{1}(D),D),\ldots,\mathcal{M}_{T}(\mathcal{M}_{1}(D),\ldots,\mathcal{M}_{T-1}(D),D)\big{)}.$
The results of (Balle and Wang,, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al., (2022).
Lemma 4 (Dong et al., 2022).
Consider an adaptive composition of $T$ Gaussian mechanisms, each with $L_{2}$-sensitivity $\Delta$ and noise scale parameter $\sigma$. The adaptive composition is $(\varepsilon,\delta)$-DP for
$\delta(\varepsilon)=\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}+\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right)-e^{\varepsilon}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}-\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right).$
2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
In our results, the network DP guarantees become those of a projected Gaussian mechanism of the form
$\mathcal{M}(D)=f(D)+AZ,$ (2.1)
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$. To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$.
We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan,, 2013).
Definition 5 (Compact SVD and Moore–Penrose Pseudoinverse).
Let $A\in\mathbb{R}^{m\times n}$ be a matrix of rank $r$. The compact SVD (also known as the economy SVD) of $A$ is given by:
$A=U_{r}\Sigma_{r}V_{r}^{\top}$
where $U_{r}\in\mathbb{R}^{m\times r}$ and $V_{r}\in\mathbb{R}^{n\times r}$ contain the left and right singular vectors corresponding to the non-zero singular values, respectively, and $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the non-zero singular values $\sigma_{1}\geq\dots\geq\sigma_{r}>0$. The Moore–Penrose pseudoinverse of $A$, denoted $A^{+}$, is defined via compact SVD as
$A^{+}=V_{r}\Sigma_{r}^{+}U_{r}^{\top}$
where $\Sigma_{r}^{+}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with entries $\frac{1}{\sigma_{1}},\dots,\frac{1}{\sigma_{r}}$, i.e., $\Sigma_{r}^{+}=\mathrm{diag}\left(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}},\dots,\frac{1}{\sigma_{r}}\right).$
To analyze the gossip averaging algorithms, the following technical result will play a central role.
Lemma 6.
Let $\sigma>0$ and $A\in\mathbb{R}^{m\times n}$. Suppose $\mathcal{M}(D)$ is a projected Gaussian mechanism of the form given in Eq. (2.1). Let $D$ and $D^{\prime}$ be two datasets such that $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, where $\mathrm{Range(A)}$ denotes the subspace spanned by the columns of $A$. Then, for all $\alpha\geq 0$,
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{\alpha}\left(\mathcal{N}\big{(}\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2},\sigma^{2}\big{)}\,||\,\mathcal{N}(0,\sigma^{2})\right).$
i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
2.3 Decentralized Learning and Network DP
We consider a connected graph $G=(V,E)$, with $V$ and $E$ as the nodes and edges of graph, respectively. The graph structure is encoded into an adjacency matrix $A\in\{1,0\}^{n\times n}$, with $A_{ij}=1$ if nodes $i$ and $j$ are connected by an edge and 0 otherwise. We denote the nodes via indices, i.e., the $n$ nodes of the graph are denoted via $[n]=\{1,\ldots,n\}$. The time is divided into discrete intervals. In each round $t\in[T]$, $T\in\mathbb{Z}_{+}$, each node $i$ can exchange a message with its neighboring nodes $N_{i}=\{j|(i,j)\in E\}$. We denote the closed neighborhood of node $i$ by $\bar{N}_{i}=N_{i}\cup\{i\}$.
The total dataset $D$ is partitioned into $n$ disjoint shards with each node $i$ holding a local dataset $D_{i}$. In decentralized learning the goal is to learn a model $\theta$ that minimizes the loss function $f(\theta)=\sum\nolimits_{i=1}^{n}f_{i}(\theta)$, where $f_{i}(\theta)$ denotes the empirical loss of node $i$, $i\in[n]$, and $\theta\in\mathbb{R}^{d}$ denotes the model parameters. At each round $t\in[T]$, the global state variable is denoted $\theta_{t}\in\mathbb{R}^{n}$ with node-wise variables $[\theta_{t}]_{i}$, $i\in[n]$. In our analysis, we consider univariate node-wise variables. Our results are generalizable to higher dimensions via tensorization using Kronecker products.
DP requires defining a neighborhood relation for datasets. In our experiments we adopt a record-level relation which describes the protection for an individual data element in a dataset $D_{i}$ of a given node $i\in[n]$, however we note that our results are applicable to any neighborhood relation. Formally, $D=\cup_{i\in[n]}D_{i}$ and $D=\cup_{i\in[n]}D_{i}$ are adjacent datasets, denoted $D\sim D^{\prime}$, if there exists $i\in[n]$ such that only $D_{i}$ and $D_{i}^{\prime}$ differ. We use $D\sim_{i}D^{\prime}$ to denote that $D$ and $D^{\prime}$ differ only in the data of node $i$.
We define a decentralized mechanism $\mathcal{M}$ as a randomized function that takes as input the total dataset $D=\cup_{i\in[n]}D_{i}$ and outputs all the shared messages. What is seen by whom, is represented by the view. We denote by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})$ the part of $\mathcal{M}(D)$ visible to the set of nodes $\mathcal{A}$. For a set of nodes $\mathcal{A}=\{i_{1},\ldots,i_{\left|\mathcal{A}\right|}\}$ we define the selector matrix $S(\mathcal{A})\in\mathbb{R}^{\left|\mathcal{A}\right|\times n}$ as $S(\mathcal{A})=\sum_{k=1}^{\left|\mathcal{A}\right|}e_{k}e_{i_{k}}$, where $e_{k}$ is the $k$th standard basis vector of $\mathbb{R}^{\left|\mathcal{A}\right|}$ and $e_{i_{k}}$ the $i_{k}$th standard basis vector of $\mathbb{R}^{n}$.
We say that the node $j$ is $(\varepsilon,\delta)$-DP from point of view of a subset of nodes $\mathcal{A}\subset[n]$, if $D\simeq_{j}D^{\prime}$ and $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\simeq_{(\varepsilon,\delta)}\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$ We also assume that nodes are honest but curious: they adhere to the protocol, but may attempt to infer additional information from the messages they observe. Similarly to (Cyffers et al.,, 2022), we first assume that the noise injected by the observing nodes contributes to the DP guarantees. In Section 3.5, we show how to remove this assumption.
[…]
3 Privacy Analysis
3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas,, 2005). In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
$\theta_{t+1}$ $=A_{t}\theta_{t}+B_{t}u_{t},$
$y_{t}$ $=C_{t}\theta_{t},$
where at each time step $t$, $\theta_{t}$ represents the state vector, the matrix $A_{t}$ the state transition matrix, $B_{t}$ the input matrix which describes how the control input $u_{t}$ at time step $t$ affects the state, and the observation $y_{t}$ is related to the state vector $\theta_{t}$ by the observation matrix $C$.
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch. 4 Antoulas,, 2005). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
This state-space dynamics perspective can be used to analyze different threat models for both non-adaptive and adaptive compositions. To summarize our results, when measuring the DP guarantees for the data of node $j$, the coefficient matrices $(A,B,C)$ are given as follows.
  1. 1.
    Non-adaptive analysis with secure summation from the view of node $i$:
    $(A,B,C)=(W,W,e^{\top}_{i}),$
    where $e_{i}$ denotes the $i$th standard basis (one-hot) vector in $\mathbb{R}^{n}$.
  2. 2.
    Non-adaptive analysis with secure summation and a set of colluding nodes $\mathcal{C}=\{i_{1},\ldots,i_{\left|\mathcal{C}\right|}\}$:
    $(A,B,C)=\big{(}W,W,S(\mathcal{C})\big{)}$
    with the selector matrix $S(\mathcal{C})$.
  3. 3.
    Non-adaptive and adaptive analysis without secure summation:
    $(A,B,C)=(W,I_{n},e_{j}).$
  4. 4.
    Adaptive analysis with secure summation:
    $(A,B,C)=\big{(}W,W,S(\bar{N}_{j})\big{)}$
    with the selector matrix $S(\bar{N}_{j})$.
We next derive these matrices and the corresponding DP guarantees one by one.
3.2 Analysis of Non-Adaptive Compositions with Secure Summation
We first show how to analyze DP gossip averaging with secure summation protocols for non-adaptive node-wise functions of data. To this end, without loss of generality, consider the task of globally averaging node-wise streams of data. I.e., each node $i$, $i\in[n]$, will have a stream $x_{0}^{i},x_{1}^{i},\dots$. Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round. For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}.$ Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
$\theta_{t+1}=W(\theta_{t}+x_{t}+u_{t}),$ (3.1)
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. From the recursion (3.1) it follows that from the point-of-view of node $i$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
$\vdots$
$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
From this representation we get the following.
Lemma 1.
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality). Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
$\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$ (3.2)
Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 1 directly gives the following result.
Theorem 2.
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2). Denote $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $H_{T}^{+}$ denotes the Moore–Penrose pseudoinverse of the matrix $H_{T}$. Then, from the point-of-view of a node $i$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
3.3 Analysis of Adaptive Compositions with Secure Summation
When analyzing adaptive compositions, the gossip averaging algorithm can be written as
$\theta_{t+1}=W(\theta_{t}+x_{t}(\theta_{t})+u_{t}),$ (3.3)
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $\theta_{t}$ denote the global state of the average estimate at round $t$. Naturally, each element of $x_{t}$ depends only on the corresponding element of $\theta_{t}$, i.e., for any $i\in[n]$, $[x_{t}(\theta_{t})]_{i}=[x_{t}([\theta_{t}]_{i})]_{i}$.
Central to our analysis is the observation that in case $D\simeq_{j}D^{\prime}$, in case of both $D$ and $D^{\prime}$, the view of each node $i$ in the graph is post-processing of the view of $\bar{N_{j}}$, the closed neighborhood of the node $j$, i.e., for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained from post-processing of the views $\mathrm{View}_{\mathcal{M}(D)}\big{(}\bar{N_{j}}\big{)}$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}\big{(}\bar{N_{j}}\big{)}$, respectively. As we show, this also allows analyzing adaptive compositions.
Notice that the selector matrix $S(\bar{N}_{j})$ selects the rows of $W$ corresponding to the indices in the set of closed neighborhood $\bar{N}_{j}$. More specifically, if $\left|\bar{N}_{j}\right|=d_{j}$ and $\bar{N}_{j}=\{i_{1},\ldots,i_{d_{j}}\}$, then
$S(\bar{N}_{j})W=\begin{bmatrix}e_{i_{1}}^{\top}W\\ \vdots\\ e_{i_{d_{j}}}^{\top}W\end{bmatrix}\in\mathbb{R}^{d_{j}\times n}.$
Denoting $S_{j}:=S(\bar{N}_{j})$, from the point-of-view of $\bar{N}_{j}$, what is observed is then given by $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$ (3.4)
$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
$\vdots$
$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})\cdots+W(x_{T-1}+u_{T-1})\big{)}.$
With a similar derivation as in the case of non-adaptive compositions, this representation gives the following result.
Theorem 3.
Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$. Denote
$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
We now turn to the analysis of node-level differential privacy guarantees in the absence of secure summation, where nodes can observe the plain messages sent by their neighbors. In this setting, we analyze both non-adaptive and adaptive compositions simultaneously. However, we note that for the non-adaptive case, our techniques can yield tighter bounds by incorporating the distance between the observing node and the node whose privacy is being evaluated.
3.4 Gossip Averaging without Secure Summation
In case the nodes are allowed to see the plain locally perturbed messages of their neighboring nodes, the analysis changes. Now, crucial observation is that the view of any node in the graph is post-processing of the view of the differing node only. I.e., in case $D\simeq_{j}D^{\prime}$, then for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained by applying the same $j$th node-independent post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{j\})$, respectively.
To analyze the view $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$, instead of the iteration (3.3), we have
$\mathrm{View}_{\mathcal{M}(D)}(\{j\})=\begin{bmatrix}e_{j}^{\top}\theta_{1}\\ \vdots\\ e_{j}^{\top}\theta_{T}\end{bmatrix}$
where
$\theta_{t+1}=W\theta_{t}+x_{t}+u_{t},\quad\theta_{0}=0,$ (3.5)
Writing out the recursion as in Eq. (3.4) and following the same steps, gives the following result.
Theorem 4.
Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
3.5 Accounting for Nodes’ Knowledge of Injected Noise
As stated in Section 2, we have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes. In practice, this likely has a small effect on the the DP guarantees. However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees.
For example, in case of non-adaptive compositions and secure summation, instead of Thm. 2, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big{(}T\cdot(n-1)\big{)}$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. Similar correction can be carried out also in all the other cases. We do this correction in all our experiments.
4 Experiments
Illustration of the DP Bounds for Non-Adaptive Compositions. We first consider numerical evaluation of the DP bounds applicable for non-adaptive compositions (Thm. 2 and Thm. 4).
First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$. We take the gossip matrix $W$ to be a doubly stochastic matrix obtained via a max degree normalization, where we first scale the off-diagonal elements as $W_{ij}=A_{ij}/\max\{d_{i},d_{j}\}$, where $A$ denotes the symmetric adjacency matrix and $d_{i}$ the degree of node $i$, $i\in[n]$. We then set the diagonal elements as $W_{ii}=1-\sum_{j,i\neq j}W_{ij}$. The left figure of Fig. 1 illustrates that in case of the secure summation, the squared sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}$ for a randomly chosen $j$ grows linearly w.r.t. $T$, and in particular, approaches $\frac{T}{n}$ which corresponds to the sensitivity in the case of centralized aggregation.
Second, consider the Facebook Social Circle dataset (Leskovec and Mcauley,, 2012) available in the Stanford Large Network Dataset Collection (Leskovec and Sosič,, 2016), which describes an directed graph of $n=4036$ nodes connected via binary edge weights. We consider an undirected version of this graph that we obtain via symmetrization of the adjacency matrix. We obtain a doubly stochastic gossip matrix $W$ using the max degree normalization similarly to the synthetic case. The right figure of Fig. 1 illustrates the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ for observing the node $j$ from node $i$, where $j$ is the node index 300 (an arbitrary node) and $i$ is either one of the first two indices in the neighborhood of node $j$ or one of the first two indices outside of the neighborhood of node $j$.
Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: Facebook Social Circle dataset (Leskovec and Mcauley,, 2012), and the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ as a function of $T$. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not. In case of the synthetic Erdős–Rényi graph, the scaled sensitivity approaches $\frac{1}{n}$ which corresponds to the scaled sensitivity in the centralized case using secure aggregation.
Illustration of the DP Bounds for Adaptive Compositions. We also consider comparing the bounds in case of adaptive compositions (Thm. 3 and Thm. 4), which corresponds to gossip averaging based ML model training via private gradient descent, for example. We consciously exclude the composition bound of (Cyffers et al.,, 2022, Thm. 6) in our comparisons. Their bound exhibits an $O(T^{2})$ dependency in the Rényi DP parameters and becomes even looser than the standard LDP bound in our setting (when the number of Muffliato averaging steps in their algorithm is $K=1$). We compute average the $\varepsilon$-values of the pair-wise DP guarantees, i.e., the average of the DP-$\varepsilon$’s between all pairs $(i,j)$ over the whole graph, when $\delta=10^{-5}$. Interestingly, the bound we obtain using Thm. 4 for the averaging without secure summation becomes tighter than the bound for the averaging with secure summation (Thm. 3) as $T$ grows. This suggests that the bound of Thm. 3 can be improved further. Notably, both bounds represent clear improvements over the LDP bound.
Figure 2: DP amplification for adaptive compositions over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=50$ and $p=0.2$.
Decentralized Learning for Logistic Regression. We distribute the MNIST (LeCun,, 1998) dataset IID to 100 nodes and each node is minimizing the cross-entropy loss for the logistic regression model using gradient descent. The nodes are connected based on an adjacency matrix $A\in\{0,1\}^{n\times n}$ that corresponds to a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$. We compute the average DP-$\varepsilon$ values of all pairs $(i,j)$ (DP of node $j$ from the point of view of node $i$) over the whole graph, when $\delta=10^{-5}$. The messages of the nodes are locally DP updated models and nodes carry out gossip averaging over the closed neighborhoods, i.e., the row-stochastic gossip matrix $W=\Lambda^{-1}(A+I_{n})$, where $\Lambda$ is a diagonal matrix with $\Lambda_{ii}=\left|\bar{N_{i}}\right|^{-1}$, $i\in[n]$.
Figure 3: Decentralized learning for logistic regression on MNIST dataset over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$. The gossip matrix is the row-stochastic matrix corresponding to averaging over the closed neighborhood of each node.
5 Conclusions and Outlook
In this work, we have given the first privacy amplification results for DP gossip averaging algorithms on graphs that are able to accurately capture the privacy amplification arising from all the noise injected in the system. As a future work, it will be interesting to see, whether the strong results for non-adaptive compositions that use secure summation can be translated to bounds for adaptive compositions. From computational perspective, a necessary future task will be to speed up linear algebraic subroutines for evaluating the total sensitivities for large and possibly sparse graphs. Also, the proposed framework is likely applicable to other decentralized optimization methods exhibiting a similar linear structure as the gossip averaging considered here.
Appendix A Proof of Lemma 6
Lemma 1.
Let $\sigma>0$ and $A\in\mathbb{R}^{m\times n}$. Suppose $\mathcal{M}(D)$ is a projected Gaussian mechanism of the form given in Eq. (2.1). Let $D$ and $D^{\prime}$ be two datasets such that $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, where $\mathrm{Range(A)}$ denotes the subspace spanned by the columns of $A$. Then, for all $\alpha\geq 0$,
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{\alpha}\left(\mathcal{N}\big{(}\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2},\sigma^{2}\big{)}\,||\,\mathcal{N}(0,\sigma^{2})\right).$
i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
Proof.
Recall, the projected Gaussian mechanism is give as
$\mathcal{M}(D)=f(D)+AZ,$
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$.
Denote the compact SVD of $A$ as $A=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $A$. Since the columns of $U_{r}$ give a basis for the subspace $\mathrm{Range(A)}$ and $U_{r}$ has orthonormal columns, $U_{r}U_{r}^{\top}$ gives a projector onto $\mathrm{Range(A)}$. Since $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, it holds that
$f(D)-f(D^{\prime})=U_{r}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}.$ (A.1)
By using the compact SVD of $A$, we have that for all $\alpha\geq 0$:
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)$ $=H_{\alpha}\left(f(D)+AZ\,||\,f(D^{\prime})+AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+AZ\,||\,AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\Sigma_{r}V_{r}^{\top}Z\,||\,\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+V_{r}^{\top}Z\,||\,V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
where $\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2}I_{r})$ and where before the third last equality we have carried out multiplication from the left by $U_{r}^{\top}$ and before the second last equality we have carried out multiplication from the left by $\Sigma_{r}^{-1}$. The last step follows from the fact that $V_{r}$ has orthonormal columns. We see that for all $\varepsilon\in\mathbb{R}$,
$H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
where the right-hand side gives the tight $\delta(\varepsilon)$ for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}\right\rVert_{2}$ and noise scale $\sigma$. Further, since $V_{r}$ has orthonormal columns,
$\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}\right\rVert_{2}$ $=\left\lVert V_{r}\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}\right\rVert_{2}$
$=\left\lVert A^{+}\big{(}f(D)-f(D^{\prime})\big{)}\right\rVert_{2}$
and the claim follows. ∎
Appendix B Proof of Lemma 1
Lemma 1.
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality). Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
$\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$
Proof.
We now need to bound the hockey-stick divergence
$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right).$
Recall that from the recursion (3.1) it follows that
$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix},$
where
$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
$\vdots$
$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
I.e.,
$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
$\widehat{x}_{T}=\begin{bmatrix}x_{0}\\ x_{1}\\ \vdots\\ x_{T-1}\end{bmatrix}.$
similarly,
$\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})=H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
$\widehat{x}_{T}^{\prime}=\begin{bmatrix}x_{0}^{\prime}\\ x_{1}^{\prime}\\ \vdots\\ x_{T-1}^{\prime}\end{bmatrix}$
for some $\widehat{x}_{T}^{\prime}\in\mathbb{R}^{T\cdot n}$ such that
$\widehat{x}_{T}-\widehat{x}_{T}^{\prime}=\begin{bmatrix}e_{j}\Delta_{1}\\ e_{j}\Delta_{2}\\ \vdots\\ e_{j}\Delta_{n}\end{bmatrix},$
where $e_{j}$ denotes the $j$th standard basis vector in $\mathbb{R}^{n}$ and $\left|\Delta_{i}\right|\leq 1$ for all $i\in[T]$.
From the translational invariance of the hockey-stick divergence, we have
$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right)$ $=H_{\alpha}\left(H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(H_{T}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right).$
By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
$H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(\Sigma_{r}V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+V_{r}^{\top}\widetilde{u}_{T}\,||\,V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+\widehat{u}\,||\,\widehat{u}\right)$
where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$.
Denoting $\Delta x:=\widehat{x}_{T}-\widehat{x}_{T}^{\prime}$, we see that $-1\leq(\Delta x)_{i}\leq 1$ for all $i\in[T]$. For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain $[0,1]^{T}$. We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in the convex and compact domain $[0,1]^{T}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$. Thus, the optimum is found from somewhere at the boundary, i.e., it holds that $\left|(\Delta x)_{i}\right|=1$ for all $i\in[T]$. Moreover, since the rows of $H_{T}$ are vectors with non-negative entries, $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is maximized when $\Delta x=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$. ∎
Appendix C Proof of Thm. 3
Theorem 1.
Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$. Denote
$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
Proof.
Suppose $D\simeq_{j}D^{\prime}$. Denote $S_{j}=S(\bar{N}_{j})$. We need to carry out the analysis for the view $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$
$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
$\vdots$
$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})+\cdots+W(x_{T-1}+u_{T-1})\big{)},$
where $[x_{i}]_{k}$ depends on the state $[\theta_{i}]_{k}$ (or could depend also on $[\theta_{i-1}]_{k},[\theta_{i-2}]_{k},\ldots$).
The only source of randomness in $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ is that of $u_{0},\ldots,u_{T-1}$. To carry out the DP analysis, we "roll out" by starting the integration of the hockey-stick divergence from the last noise vector $u_{T-1}$. This is similar to the composition analysis using RDP (Mironov,, 2017) and using dominating pairs of distributions (Zhu et al.,, 2022). In particular, the proof is similar to the proof of (Thm. 10, Zhu et al.,, 2022).
The general case can be illustrated using with the case $T=2$. When $T=2$, we need to analyze the mechanism
$\mathcal{M}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},$
where $[x_{1}]_{k}$ depends on the state $[\theta_{1}]_{k}$ for all $k\in[n]$.
The essential observation here is that when $D\simeq_{j}D^{\prime}$, the the states of the nodes outside of the closed neighborhood $\bar{N}_{j}$ are obtained by applying the same post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\bar{N}_{j})$. More specifically, given the view of the nodes inside the closed neighborhood $\bar{N}_{j}$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the nodes outside of $\bar{N}_{j}$, the vectors $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution, since they consist of linear combinations of the entries of $y_{1}$ and of entries of $\theta_{1}$ outside of $\bar{N}_{j}$ which are distributed similarly in case of $D$ and $D^{\prime}$.
Therefore, when starting to integrate with respect to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and we can also fix all the terms in $Wx_{1}$ that are evaluated using the states of the nodes outside of $\bar{N}_{j}$. What remains in $Wx_{1}$ are the terms evaluated using $y_{1}$, and thus we can also carry out an adaptive analysis. We then roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
where
$\mathcal{M}_{1}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}$
and
$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+Wu_{1}\big{)}\end{bmatrix}.$
Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
where
$\mathcal{M}_{2}(D)=\begin{bmatrix}S_{j}\big{(}W(e_{j}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(e_{j}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}\big{(}W^{2}+W\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
and
$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}Wu_{0}\big{)}\\ S_{j}\big{(}W^{2}u_{0}+Wu_{1}\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
In case $T=3$, we can use the same reasoning, and similarly, given the observations $y_{1}$ and $y_{2}$, we can freeze the terms $S_{j}\big{(}W^{3}(x_{0}+u_{0})$ and $S_{j}W^{2}(x_{1}+u_{1})$ when integrating w.r.t. $u_{2}$. The general case also follows from this reasoning.
Appendix D Proof of Thm. 4
Theorem 1.
Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
Proof.
The proof goes similarly as the proof of Thm. 3 and can be illustrated using the case $T=2$.
Given the view of the node $j$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the rest of the nodes, the vectors $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution.
Therefore, when starting to integrate w.r.t. to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
where
$\mathcal{M}_{1}(D)=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}$
and
$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+u_{1}\big{)}\end{bmatrix}$
Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
where
$\mathcal{M}_{2}(D)=\begin{bmatrix}e_{j}^{\top}(e_{j}+u_{0})\\ e_{j}^{\top}\big{(}W(e_{j}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}1\\ e_{j}^{\top}\big{(}W+I\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
and
$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}Wu_{0}\\ e_{j}^{\top}\big{(}Wu_{0}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
The general case follows from the same reasoning, similarly as in the proof of Thm. 3.
Appendix E Results for Time-Varying Graphs
We see from the proofs that all the results can be straightforwardly generalized to time-varying gossip matrices $W_{t}$, $t\in[T]$. For example, in case of the non-adaptive compositions with secure summation, we would have Theorem 1 with
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}W_{0}e_{j}\\ e_{i}^{\top}(W_{1}\cdot W_{0}+W_{0})e_{j}\\ \vdots\\ e_{i}^{\top}\left(\prod\limits_{i=0}^{T-1}W_{i}+e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}+\cdots+W_{0}\right)e_{j}\end{bmatrix}$
and
$H_{T}:=\begin{bmatrix}e_{i}^{\top}W_{0}&0&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}&e_{i}^{\top}W_{1}&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}\cdot W_{2}&e_{i}^{\top}W_{1}\cdot W_{2}&e_{i}^{\top}W_{2}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}\prod\limits_{i=0}^{T-1}W_{i}&e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}&\dots&e_{i}^{\top}W_{T-2}\cdot W_{T-1}&e_{i}^{\top}W_{T-1}\end{bmatrix}$
The vector $\widetilde{x}_{T}$ and the matrix $H_{T}$ of Theorems 3 and 4 can be modified analogously.
Appendix F Details for Experiments
F.1 Additional Details
  • For each model, both when using the DP Decentralized Averaging and DP-FedAvg, we optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$.
  • In case of the gossip averaging, the test accuracies are evaluated on a model of a randomly chosen node.
  • The experiments on the logistic regression including all the learning rate tuning took around 48 hours using 8 RTX 3080 GPUs.
F.2 Baseline Method for Experiments
As a baseline algorithm, we consider the full batch version of the DP Federated Averaging algorithm depicted in Alg. 11. (Ponomareva et al.,, 2023; McMahan et al.,, 2017)
Algorithm 1 Differentially Private Federated Averaging (DP-FedAvg) with Record-Level DP
1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise parameter $\sigma$, initial global model $w^{0}$
2:  for each round $t=0$ to $T-1$ do
3:     for each client $i$ in parallel do
4:        Initialize local model: $w_{i}^{t}\leftarrow w^{t}$
5:        Perform DP gradient descent step on $D_{i}$, with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma/\sqrt{n}$ obtaining updated model $w_{i}^{t}$
6:        Compute model update: $\Delta_{i}^{t}\leftarrow w_{i}^{t}-w^{t}$
7:     end for
8:     Server aggregates updates:
$\tilde{\Delta}^{t}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}^{t}.$
9:     Update global model:
$w^{t+1}\leftarrow w^{t}+\tilde{\Delta}^{t}$
and distribute it to clients.
10:  end for
11:  Output: Final global model $w^{T}$
F.3 Gossip Averaging Algorithm
In the experiments, we consider the full decentralized DP gradient descent shown in Alg. 10.
Algorithm 2 Differentially Private Gossip Averaging (DP-GossipAvg)
1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise multiplier $\sigma$, gossip matrix $W\in\mathbb{R}^{n\times n}$, initial local models $\{w_{i}^{0}\}_{i=1}^{n}$
2:  for each round $t=0$ to $T-1$ do
3:     for each client $i$ in parallel do
4:        Perform local DP gradient descent on $D_{i}$ with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma$ to obtain updated local model $w_{i}^{t}$
5:     end for
6:     for each client $i$ in parallel do
7:        Gossip averaging:
$w_{i}^{t+1}\leftarrow\sum_{j=1}^{n}W_{ij}w_{j}^{t}$
8:     end for
9:  end for
10:  Output: Final local models $\{w_{i}^{T}\}_{i=1}^{n}$

### v2

@@
 Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance.
-However, achieving differential privacy (DP) in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes.
-In this work, we present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, both with and without secure summation over each node’s direct neighbors.
-Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage across these scenarios.
-This framework significantly improves upon prior analyses, for example, reducing the Rényi DP parameter growth from $O(T^{2})$ to $O(T)$, where $T$ is the number of training rounds.
-We validate our analysis with numerical results demonstrating superior DP bounds compared to existing approaches.
-We further illustrate our analysis with a logistic regression experiment on MNIST image classification in a fully decentralized setting, demonstrating utility comparable to central aggregation methods.
+However, achieving differential privacy (DP) guarantees in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes.
+We present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph and especially consider the averaging over nearest neighbors with secure summation and individual node-wise views.
+Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes across different scenarios.
+In case the gossip averaging happens via secure summation, we show that the Rényi DP parameter growth is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
 1 Introduction
@@
 Fully decentralized gradient-based optimization methods are typically distributed variants of gradient descent and can be broadly classified into two categories.
-In random walk-based methods (Lopes and Sayed,, 2007; Johansson et al.,, 2009; Mao et al.,, 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix.
+In random walk-based methods (rw1; rw2; rw3), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix.
 The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor.
@@
 However, such sequential update dynamics inherently limits parallelism and scalability.
-In gossip averaging-based methods (Boyd et al.,, 2006; Dimakis et al.,, 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states.
+In gossip averaging-based methods (gossip2; gossip1), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states.
 Each node updates its parameters by averaging over the received messages.
@@
 DP in Centralized FL.
-In centralized FL (Kairouz et al., 2021b, ), distributed DP (Ullah et al.,, 2023; Hartmann and Kairouz,, 2023; Kairouz et al., 2021a, ) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al.,, 2019; Erlingsson et al.,, 2019; Bell et al.,, 2020).
+In centralized FL (FLbook), distributed DP (UllahCKO23; google_flblog_2023; FLDP3) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (truex2019hybrid; Erlingsson2019; bell2020secure).
 This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
 DP in Decentralized Optimization.
-The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al.,, 2025; Zhu et al.,, 2019; Mrini et al.,, 2024; Pasquini et al.,, 2023).
-Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al.,, 2015; Bellet et al.,, 2018; Xu et al.,, 2022), relying on local DP guarantees.
-This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al.,, 2012).
-To address this, recent works propose pairwise network DP (Cyffers et al.,, 2024, 2022; Cyffers and Bellet,, 2022), where privacy loss is analyzed between pairs of nodes.
-These relaxations yield better utility while accounting for the structure of decentralized communication.
-However, for the practical case of gossip averaging, there does not yet exist a satisfactory composition analysis for the pairwise network DP guarantees.
-To the best of our knowledge, the best bounds in this case are the Rényi differential privacy (RDP) bounds by Cyffers et al., (2022) which exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of training iterations, making them unsuitable for practical ML model training with reasonable privacy guarantees.
+The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (dlattack4; dlattack3; dlattack1; dlattack2).
+Early decentralized DP methods achieve privacy by locally perturbing gradients or models (HMV2015; BGTT18; XZW2022), relying on local DP guarantees.
+This often imposes a poor utility-privacy tradeoff compared to centralized DP (ChanSS12).
+To address this, recent works propose pair-wise network DP (networkdp3; networkdp2; networkdp1), where privacy loss is analyzed between pairs of nodes.
+These relaxations give better utility while accounting for the structure of decentralized communication.
+networkdp2 give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees.
+Recently, bellet2025unified considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (kairouz2021practical; pillutla2025correlated) which leads to considerably more accurate privacy analysis than that of networkdp2.
+Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
 Our Approach.
-In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows empirically as $O(\sqrt{T})$, resulting in $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses.
+In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses.
 This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs).
 This also provides tools to analyze the network DP guarantees in different threat scenarios.
-In addition to considering the threat model of (Cyffers et al.,, 2022), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
+In addition to considering the threat model of (networkdp2), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
 Secure Aggregation in Decentralized Learning.
 While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored.
-In current gossip-based systems (Cyffers et al.,, 2022), each node receives messages directly from its neighbors and can, in principle, view their full content.
-As we show, summation over a node’s neighborhood can yield a meaningful privacy amplification.
-Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al.,, 2018; Lian et al.,, 2018; Jeon et al.,, 2021).
-Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al.,, 2022; Pereira et al.,, 2024; Biswas et al.,, 2024).
-The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al.,, 2024).
+In current gossip-based systems (networkdp2), each node receives messages directly from its neighbors and can, in principle, view their full content.
+As we show, summation over a node’s neighborhood can give a meaningful privacy amplification.
+Several studies have proposed integrating MPC primitives into decentralized learning workflows (jayaraman2018distributed; lian2018asynchronous; jeon2021privacy).
+Concrete protocols for decentralized secure aggregation have been developed in recent works (sabater2022gopa; dlsa1; biswas2024).
+The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (nicolas2024differentially).
 Our Contributions:
-  • We develop a novel analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms.
+  • We provide an analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms.
 Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
-  • We demonstrate that the DP guarantees, for both non-adaptive and adaptive compositions-with or without secure summation-can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
-  • Our empirical results show that the squared sensitivity scales as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on Rényi differential privacy (RDP).
-  • We validate our findings on a decentralized logistic regression task using the MNIST dataset, showing privacy-utility trade-offs comparable to those of central FL.
+  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
+  • We show that in case secure summation is used to carry averaging over the nearest neighbors, the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (networkdp2).
 2 Background
@@
 We say that a mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable.
-Definition 1 (Dwork et al., 2006).
+Definition 1 (dwork_et_al_2006).
 Let $\varepsilon\geq 0$ and $\delta\in[0,1]$.
@@
 With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions.
-For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $[z]_{+}=\max\{0,z\}$.
+For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $z]_{+}=\max\{0,z\}$.
 The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
-Lemma 2 (Balle et al., 2018).
+Lemma 2 (balle2018subsampling).
 A mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if and only if, $\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\delta$.
@@
 $\mathcal{M}^{(T)}(D)=\big(\mathcal{M}_{1}(D),\mathcal{M}_{2}(\mathcal{M}_{1}(D),D),\ldots,\mathcal{M}_{T}(\mathcal{M}_{1}(D),\ldots,\mathcal{M}_{T-1}(D),D)\big).$
-The results of (Balle and Wang,, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism.
-Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al., (2022).
-Lemma 4 (Dong et al., 2022).
+The results of (balle2018gauss) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism.
+Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by dong2022gaussian.
+Lemma 4 (dong2022gaussian).
 Consider an adaptive composition of $T$ Gaussian mechanisms, each with $L_{2}$-sensitivity $\Delta$ and noise scale parameter $\sigma$. The adaptive composition is $(\varepsilon,\delta)$-DP for
@@
 To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$.
-We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan,, 2013).
+We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (golub2013matrix).
 Definition 5 (Compact SVD and Moore–Penrose Pseudoinverse).
@@
 i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
-2.3 Decentralized Learning and Network DP
-We consider a connected graph $G=(V,E)$, with $V$ and $E$ as the nodes and edges of graph, respectively.
-The graph structure is encoded into an adjacency matrix $A\in\{1,0\}^{n\times n}$, with $A_{ij}=1$ if nodes $i$ and $j$ are connected by an edge and 0 otherwise.
-We denote the nodes via indices, i.e., the $n$ nodes of the graph are denoted via $[n]=\{1,\ldots,n\}$.
-The time is divided into discrete intervals.
-In each round $t\in[T]$, $T\in\mathbb{Z}_{+}$, each node $i$ can exchange a message with its neighboring nodes $N_{i}=\{j|(i,j)\in E\}$.
-We denote the closed neighborhood of node $i$ by $\bar{N}_{i}=N_{i}\cup\{i\}$.
-The total dataset $D$ is partitioned into $n$ disjoint shards with each node $i$ holding a local dataset $D_{i}$.
-In decentralized learning the goal is to learn a model $\theta$ that minimizes the loss function $f(\theta)=\sum\nolimits_{i=1}^{n}f_{i}(\theta)$, where $f_{i}(\theta)$ denotes the empirical loss of node $i$, $i\in[n]$, and $\theta\in\mathbb{R}^{d}$ denotes the model parameters.
-At each round $t\in[T]$, the global state variable is denoted $\theta_{t}\in\mathbb{R}^{n}$ with node-wise variables $[\theta_{t}]_{i}$, $i\in[n]$.
-In our analysis, we consider univariate node-wise variables.
-Our results are generalizable to higher dimensions via tensorization using Kronecker products.
-DP requires defining a neighborhood relation for datasets.
-In our experiments we adopt a record-level relation which describes the protection for an individual data element in a dataset $D_{i}$ of a given node $i\in[n]$, however we note that our results are applicable to any neighborhood relation.
-Formally, $D=\cup_{i\in[n]}D_{i}$ and $D=\cup_{i\in[n]}D_{i}$ are adjacent datasets, denoted $D\sim D^{\prime}$, if there exists $i\in[n]$ such that only $D_{i}$ and $D_{i}^{\prime}$ differ.
-We use $D\sim_{i}D^{\prime}$ to denote that $D$ and $D^{\prime}$ differ only in the data of node $i$.
-We define a decentralized mechanism $\mathcal{M}$ as a randomized function that takes as input the total dataset $D=\cup_{i\in[n]}D_{i}$ and outputs all the shared messages.
-What is seen by whom, is represented by the view.
-We denote by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})$ the part of $\mathcal{M}(D)$ visible to the set of nodes $\mathcal{A}$.
-For a set of nodes $\mathcal{A}=\{i_{1},\ldots,i_{\left|\mathcal{A}\right|}\}$ we define the selector matrix $S(\mathcal{A})\in\mathbb{R}^{\left|\mathcal{A}\right|\times n}$ as $S(\mathcal{A})=\sum_{k=1}^{\left|\mathcal{A}\right|}e_{k}e_{i_{k}}$, where $e_{k}$ is the $k$th standard basis vector of $\mathbb{R}^{\left|\mathcal{A}\right|}$ and $e_{i_{k}}$ the $i_{k}$th standard basis vector of $\mathbb{R}^{n}$.
-We say that the node $j$ is $(\varepsilon,\delta)$-DP from point of view of a subset of nodes $\mathcal{A}\subset[n]$, if $D\simeq_{j}D^{\prime}$ and $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\simeq_{(\varepsilon,\delta)}\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$ We also assume that nodes are honest but curious: they adhere to the protocol, but may attempt to infer additional information from the messages they observe.
-Similarly to (Cyffers et al.,, 2022), we first assume that the noise injected by the observing nodes contributes to the DP guarantees.
-In Section 3.5, we show how to remove this assumption.
-[…]
+2.3 Decentralized Learning and Network Differential Privacy
+We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an adjacency matrix
+$A\in\{0,1\}^{n\times n},\qquad A_{ij}=1\;\text{iff}\;(i,j)\in E.$
+We denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
+$N_{i}=\{j\in[n]:(i,j)\in E\},\qquad\bar{N}_{i}=N_{i}\cup\{i\}.$
+Local Data and Learning Objective.
+The global dataset is partitioned across nodes:
+$D=\bigcup_{i=1}^{n}D_{i},$
+where node $i$ holds the shard $D_{i}$. The goal of decentralized learning is to minimize the global loss
+$f(\theta)=\sum_{i=1}^{n}f_{i}(\theta),\qquad\theta\in\mathbb{R}^{d},$
+where $f_{i}(\theta)$ depends only on $D_{i}$.
+At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$.
+For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by bellet2025unified as described in Section 3.4.
+Dataset Adjacency.
+We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
+$D\sim_{i}D^{\prime}\quad\iff\quad D_{j}=D^{\prime}_{j}\ \text{for all }j\neq i.$
+Mechanism and Observations.
+A decentralized mechanism is a randomized mapping
+$\mathcal{M}:D\mapsto\text{(all messages exchanged during the protocol)}.$
+For a subset of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{A}$ is denoted
+$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A}).$
+To extract node-wise components, we define the selector matrix: for $\mathcal{A}=\{i_{1},\dots,i_{|\mathcal{A}|}\}$,
+$S(\mathcal{A})=\sum_{k=1}^{|\mathcal{A}|}e_{k}e_{i_{k}}^{\top}\in\mathbb{R}^{|\mathcal{A}|\times n},$
+where $e_{k}$ is the $k$th standard basis vector.
+Network Differential Privacy.
+Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
+$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\;\simeq_{(\varepsilon,\delta)}\;\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$
+We assume honest-but-curious nodes: they follow the protocol but analyze observed messages.
+As in (networkdp2), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
+Gossip-Based Averaging.
+Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
+$\theta_{t+1}=Wx_{t},$
+where $W\in\mathbb{R}^{n\times n}$ is the gossip matrix. We call $W$:
+$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},\qquad\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
+For the error analysis we also consider the definition of primitive gossip matrices.
+Definition 2.1 (Primitive Matrix).
+A matrix $W\in\mathbb{R}^{n\times n}$ is called primitive if there exists $k\in\mathbb{N}$ such that
+$W^{k}>0\quad\text{(entrywise)}.$
+Equivalently, a row-stochastic matrix is primitive if and only if the induced Markov chain is irreducible and aperiodic.
+In this case, $W$ has a unique stationary distribution $\pi>0$ with $\pi^{\top}W=\pi^{\top}$, and moreover
+$W^{t}\to\mathbf{1}\pi^{\top}\qquad\text{as }t\to\infty.$
 3 Privacy Analysis
 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
-We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas,, 2005).
+We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., antoulas2005approximation).
 In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
@@
 We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch.
-4 Antoulas,, 2005).
+4 antoulas2005approximation).
 In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
-This state-space dynamics perspective can be used to analyze different threat models for both non-adaptive and adaptive compositions.
-To summarize our results, when measuring the DP guarantees for the data of node $j$, the coefficient matrices $(A,B,C)$ are given as follows.
-  1. 1.
-    Non-adaptive analysis with secure summation from the view of node $i$:
-    $(A,B,C)=(W,W,e^{\top}_{i}),$
-    where $e_{i}$ denotes the $i$th standard basis (one-hot) vector in $\mathbb{R}^{n}$.
-  2. 2.
-    Non-adaptive analysis with secure summation and a set of colluding nodes $\mathcal{C}=\{i_{1},\ldots,i_{\left|\mathcal{C}\right|}\}$:
-    $(A,B,C)=\big{(}W,W,S(\mathcal{C})\big{)}$
-    with the selector matrix $S(\mathcal{C})$.
-  3. 3.
-    Non-adaptive and adaptive analysis without secure summation:
-    $(A,B,C)=(W,I_{n},e_{j}).$
-  4. 4.
-    Adaptive analysis with secure summation:
-    $(A,B,C)=\big{(}W,W,S(\bar{N}_{j})\big{)}$
-    with the selector matrix $S(\bar{N}_{j})$.
-We next derive these matrices and the corresponding DP guarantees one by one.
-3.2 Analysis of Non-Adaptive Compositions with Secure Summation
+3.2 Analysis of Non-Adaptive Compositions
 We first show how to analyze DP gossip averaging with secure summation protocols for non-adaptive node-wise functions of data.
@@
 Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round.
-For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}.$ Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
+For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}$.
+Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
 $\theta_{t+1}=W(\theta_{t}+x_{t}+u_{t}),$ (3.1)
 where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$.
-From the recursion (3.1) it follows that from the point-of-view of node $i$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
-$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
-$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
+From the recursion (3.1) it follows that from the view determined by the selector $S$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
+$y_{1}$ $=S(x_{0}+u_{0})$
+$y_{2}$ $=S(Wx_{0}+Wu_{0}+x_{1}+u_{1})$
+$y_{3}$ $=S(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1}+x_{2}+u_{2})$
 $\vdots$
-$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
-From this representation we get the following.
+$y_{T-1}$ $=S(W^{T-1}x_{0}+\cdots+x_{T-1}+W^{T-1}u_{0}+\cdots+u_{T-1}),$
+where $S$ is the selector matrix for $\mathcal{A}$
+From this representation we get the following for a fixed pair of neighboring streams of data.
 Lemma 7.
-Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality).
+Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$.
 Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
@@
 where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
-$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$ (3.2)
+$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$ (3.2)
 Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 7 directly gives the following result.
@@
 (3.2).
-Denote $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $H_{T}^{+}$ denotes the Moore–Penrose pseudoinverse of the matrix $H_{T}$.
-Then, from the point-of-view of a node $i$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
-3.3 Analysis of Adaptive Compositions with Secure Summation
-When analyzing adaptive compositions, the gossip averaging algorithm can be written as
-$\theta_{t+1}=W(\theta_{t}+x_{t}(\theta_{t})+u_{t}),$ (3.3)
-where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $\theta_{t}$ denote the global state of the average estimate at round $t$.
-Naturally, each element of $x_{t}$ depends only on the corresponding element of $\theta_{t}$, i.e., for any $i\in[n]$, $[x_{t}(\theta_{t})]_{i}=[x_{t}([\theta_{t}]_{i})]_{i}$.
-Central to our analysis is the observation that in case $D\simeq_{j}D^{\prime}$, in case of both $D$ and $D^{\prime}$, the view of each node $i$ in the graph is post-processing of the view of $\bar{N_{j}}$, the closed neighborhood of the node $j$, i.e., for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained from post-processing of the views $\mathrm{View}_{\mathcal{M}(D)}\big{(}\bar{N_{j}}\big{)}$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}\big{(}\bar{N_{j}}\big{)}$, respectively.
-As we show, this also allows analyzing adaptive compositions.
-Notice that the selector matrix $S(\bar{N}_{j})$ selects the rows of $W$ corresponding to the indices in the set of closed neighborhood $\bar{N}_{j}$.
-More specifically, if $\left|\bar{N}_{j}\right|=d_{j}$ and $\bar{N}_{j}=\{i_{1},\ldots,i_{d_{j}}\}$, then
-$S(\bar{N}_{j})W=\begin{bmatrix}e_{i_{1}}^{\top}W\\ \vdots\\ e_{i_{d_{j}}}^{\top}W\end{bmatrix}\in\mathbb{R}^{d_{j}\times n}.$
-Denoting $S_{j}:=S(\bar{N}_{j})$, from the point-of-view of $\bar{N}_{j}$, what is observed is then given by $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
-$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$ (3.4)
-$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
-$\vdots$
-$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})\cdots+W(x_{T-1}+u_{T-1})\big{)}.$
-With a similar derivation as in the case of non-adaptive compositions, this representation gives the following result.
-Theorem 3.
-Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$.
+Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$.
+Let $\widetilde{x}_{T}$ be of the form (3.2) and denote $c=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{T}\end{bmatrix}^{\top}\in\mathbb{R}^{T}$.
 Denote
-$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
-Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
-We now turn to the analysis of node-level differential privacy guarantees in the absence of secure summation, where nodes can observe the plain messages sent by their neighbors.
-In this setting, we analyze both non-adaptive and adaptive compositions simultaneously.
-However, we note that for the non-adaptive case, our techniques can yield tighter bounds by incorporating the distance between the observing node and the node whose privacy is being evaluated.
-3.4 Gossip Averaging without Secure Summation
-In case the nodes are allowed to see the plain locally perturbed messages of their neighboring nodes, the analysis changes.
-Now, crucial observation is that the view of any node in the graph is post-processing of the view of the differing node only.
-I.e., in case $D\simeq_{j}D^{\prime}$, then for any node $i\in[n]$, the views $\mathrm{View}_{\mathcal{M}(D)}(\{i\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})$ are obtained by applying the same $j$th node-independent post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{j\})$, respectively.
-To analyze the view $\mathrm{View}_{\mathcal{M}(D)}(\{j\})$, instead of the iteration (3.3), we have
-$\mathrm{View}_{\mathcal{M}(D)}(\{j\})=\begin{bmatrix}e_{j}^{\top}\theta_{1}\\ \vdots\\ e_{j}^{\top}\theta_{T}\end{bmatrix}$
-where
-$\theta_{t+1}=W\theta_{t}+x_{t}+u_{t},\quad\theta_{0}=0,$ (3.5)
-Writing out the recursion as in Eq. (3.4) and following the same steps, gives the following result.
-Theorem 4.
-Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
-$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
-Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
+$\Delta^{T}_{j\rightarrow i}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
+where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector.
+Then, from the point-of-view of a node $i$, the node $j$ is tightly $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
 3.3 Accounting for Nodes’ Knowledge of Injected Noise
-As stated in Section 2, we have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes.
+We have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes.
 In practice, this likely has a small effect on the the DP guarantees.
@@
 This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees.
-For example, in case of non-adaptive compositions and secure summation, instead of Thm.
-2, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big{(}T\cdot(n-1)\big{)}$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq.
+Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by allouahprivacy is scheme called SecLDP.
+For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm.
+8, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq.
 (3.2) with the $i$th column vector of each $T\times n$-block column removed.
@@
 We do this correction in all our experiments.
+3.4 Adaptive Compositions and Connection to Matrix-Mechanism Accounting
+The recent work by bellet2025unified shows that the linear systems perspective given in Theorem 8 can be equivalently analyzed via a so-called matrix mechanisms (kairouz2021practical; pillutla2025correlated).
+This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis.
+We next shortly discuss this connection.
+At a high level, both frameworks reduce an observer’s view to a linear image of a data-dependent signal plus correlated Gaussian noise.
+Different Threat Scenarios.
+The matrix mechanism formulation by bellet2025unified allows analyzing private decentralized optimization methods that have a linear structure.
+Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
+  • The LDP approach where the adversary sees all messages (bellet2018personalized). This corresponds to the setting of Theorem 8 with $S=I_{n}$.
+  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (networkdp2).
+This corresponds to Theorem 8 with a selector matrix of a finite collection of nodes.
+  • SecLDP (allouahprivacy), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees.
+This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Thmeorem 8, see Section 3.3.
+Privacy Analysis of Matrix Mechanisms.
+The matrix-mechanism analysis by bellet2025unified provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings.
+The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
+In the matrix mechanism formulation, the $d$-dimensional user-wise messages are stacked into a global matrix $G\in\mathbb{R}^{nT\times d}$ of the form
+$G=\begin{bmatrix}G_{1}\\ \vdots\\ G_{T}\end{bmatrix}$,
+where each block $G_{t}\in\mathbb{R}^{n\times d}$ corresponds to round $t$ and is given by
+$G_{t}=\begin{bmatrix}g_{t}^{1}\\ \vdots\\ g_{t}^{n}\end{bmatrix}$,
+with $g_{t}^{i}$ representing the message of node $i$ at iteration $t$.
+Thus, differences between neighboring datasets appear as differences along the rows $(g_{t}^{i})_{t=1}^{T}$ corresponding to a single user.
+The privacy analyses of different private optimization methods under varying threat scenarios are then analyzed via the following results.
+Theorem 9 (Bellet et al., 2025, Thm. 5).
+For each of the three trust models, LDP, PNDP and SecLDP, and for every linear decentralized learning algorithm, there exist matrices $A$, $B$ and $C$ such that the adversarial view can be written as
+$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (3.3)
+with $A=BC$.
+To quantify privacy, one needs to bound how much $G$ may vary under the adjacency relation while remaining observable through this linear mapping.
+Definition 3.1 (Sensitivity under participation).
+Let $B$ and $C$ be two matrices, and let $\Pi$ be a participation scheme for $G$. The sensitivity of $C$ with respect to $B$ under $\Pi$ is defined as
+$\mathrm{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\,\|\,C(G-G^{\prime})\,\|_{B^{+}B},$
+where $\|v\|_{B^{+}B}:=\sqrt{v^{\top}(B^{+}B)v}$ denotes the seminorm induced by $B^{+}B$, and $G\simeq_{\Pi}G^{\prime}$ indicates that $G$ and $G^{\prime}$ differ only along entries that belong to the same individual according to $\Pi$.
+A participation scheme $\Pi$ specifies which coordinates of the stacked update matrix $G$ may change under a neighboring-dataset relation.
+In other words, $\Pi$ encodes which entries of $G$ correspond to contributions of the same individual across time, ensuring that sensitivity is computed only over data differences permitted by the privacy definition.
+Theorem 10 (Bellet et al., 2025, Thm. 7).
+Let $\mathcal{O}_{\mathcal{A}}=AG+BZ$ be the attacker knowledge of a trust model on a linear decentralized learning algorithm, and denote by $\mathcal{M}(G)$ the corresponding mechanism.
+Let $\Pi$ be a participation scheme for $G$, and suppose $Z\sim\mathcal{N}(0,\nu^{2}I)$.
+Assume that $A$ is a column-echelon matrix and that $A=BC$ for some $C$.
+Then, if
+$\nu=\sigma\,\mathrm{sens}_{\Pi}(C;B)\qquad\text{with}\qquad\mathrm{sens}_{\Pi}(C;B)\;\leq\;\max_{\pi\in\Pi}\;\sum_{s,t\in\pi}\bigl|\bigl(C^{\top}B^{+}BC\bigr)_{s,t}\bigr|,$
+the mechanism $\mathcal{M}$ satisfies $\tfrac{1}{\sigma}$-Gaussian Differential Privacy, even when $G$ is chosen adaptively.
+Adaptive Analysis.
+Notably, bellet2025unified extend the existing adaptive analysis of matrix mechanisms (denisov2022improved) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns.
+The crucial implication is that once the representation
+$\mathcal{O}_{\mathcal{A}}=AG+BZ$
+is identified, privacy guarantees for adaptive compositions are the same as in the case of non-adaptive compositions and directly follow from a single sensitivity bound on $(B,C)$ under the participation scheme $\Pi$.
+Matrix Mechanism Formulation of Gossip Averaging.
+We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step.
+As shown in (Appendix A.4 bellet2025unified), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm.
+8.
+Equivalence of Sensitivity Bounds.
+We also see from Thm.
+10 that the sensitivity upper bound equals the upper bound given in our Thm.
+8 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
+$\left|C^{\top}B^{+}BC\right|=\left|VV^{\top}\right|$
+and the summation in the upper bound of Thm. 10 equals the upper bound given in our Thm. 8 in case the differing node participates in every round.
+4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
+We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
+We first need the following auxiliary lemma.
+Lemma 11.
+Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic. Then there exist constants $M<\infty$ and $\rho\in(0,1)$ such that for all $t\geq 1$
+$W^{t}\;=\;\Pi\;+\;R^{t},\qquad\Pi:=\mathbf{1}\,\pi^{\top},\qquad\|R^{t}\|\;\leq\;M\,\rho^{t},$
+where $\pi^{\top}W=\pi^{\top}$, $\pi>0$, $\sum_{i}\pi_{i}=1$.
+Remark 12.
+A row-stochastic primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value < 1, and as the proof of Lemma 11 shows, $1-\rho$ equals the spectral gap $\lambda_{1}-\max_{i\geq 2}\left|\lambda_{i}\right|$.
+Theorem 13.
+Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic with stationary distribution $\pi>0$, $\sum_{k=1}^{n}\pi_{k}=1$, $\pi^{\top}W=\pi^{\top}$.
+Let $H_{T}$ be the system matrix given in Lemma 7 with $S=e_{i}^{\top}$ (view of a single node, i.e., that of secure summation protocol) and define
+$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
+as in Theorem 8. Then
+$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
+Remark 14.
+Notice that if $W$ is doubly stochastic, $\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}=\frac{1}{n}$ and we asymptotically have the same sensitivity as in case of a central aggregator.
 5 Experiments
-Illustration of the DP Bounds for Non-Adaptive Compositions. We first consider numerical evaluation of the DP bounds applicable for non-adaptive compositions (Thm. 2 and Thm. 4).
+Illustration of the DP Bounds for Secure Summation Protocols. We first consider numerical evaluation of the DP bounds we obtain from Thm. 8.
 First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$.
-We take the gossip matrix $W$ to be a doubly stochastic matrix obtained via a max degree normalization, where we first scale the off-diagonal elements as $W_{ij}=A_{ij}/\max\{d_{i},d_{j}\}$, where $A$ denotes the symmetric adjacency matrix and $d_{i}$ the degree of node $i$, $i\in[n]$.
-We then set the diagonal elements as $W_{ii}=1-\sum_{j,i\neq j}W_{ij}$.
-The left figure of Fig.
-1 illustrates that in case of the secure summation, the squared sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}$ for a randomly chosen $j$ grows linearly w.r.t.
-$T$, and in particular, approaches $\frac{T}{n}$ which corresponds to the sensitivity in the case of centralized aggregation.
-Second, consider the Facebook Social Circle dataset (Leskovec and Mcauley,, 2012) available in the Stanford Large Network Dataset Collection (Leskovec and Sosič,, 2016), which describes an directed graph of $n=4036$ nodes connected via binary edge weights.
-We consider an undirected version of this graph that we obtain via symmetrization of the adjacency matrix.
-We obtain a doubly stochastic gossip matrix $W$ using the max degree normalization similarly to the synthetic case.
-The right figure of Fig.
-1 illustrates the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ for observing the node $j$ from node $i$, where $j$ is the node index 300 (an arbitrary node) and $i$ is either one of the first two indices in the neighborhood of node $j$ or one of the first two indices outside of the neighborhood of node $j$.
-Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: Facebook Social Circle dataset (Leskovec and Mcauley,, 2012), and the scaled sensitivity $(\Delta^{T}_{j\rightarrow i})^{2}/T$ as a function of $T$.
+We take the gossip matrix $W$ to be a row-stochastic matrix obtained from row-wise averaging.
+Second, we consider a network generated using a preferential-attachment model (barabasi1999emergence) with $n=200$ nodes, initialized from a fully connected core of 5 nodes.
+Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees.
+The gossip matrix $W$ is obtained from row-wise averaging.
+As Figure 1 illustrates, in case of the secure summation, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ for a randomly chosen $j$ grows asymptotically linearly w.r.t. $T$ as predicted by Thm. 13.
+Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: gossip matrix of a preferential-attachment model with $n=200$ nodes, and the scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$.
 We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
-In case of the synthetic Erdős–Rényi graph, the scaled sensitivity approaches $\frac{1}{n}$ which corresponds to the scaled sensitivity in the centralized case using secure aggregation.
-Illustration of the DP Bounds for Adaptive Compositions.
-We also consider comparing the bounds in case of adaptive compositions (Thm.
-3 and Thm.
-4), which corresponds to gossip averaging based ML model training via private gradient descent, for example.
-We consciously exclude the composition bound of (Cyffers et al.,, 2022, Thm.
-6) in our comparisons.
-Their bound exhibits an $O(T^{2})$ dependency in the Rényi DP parameters and becomes even looser than the standard LDP bound in our setting (when the number of Muffliato averaging steps in their algorithm is $K=1$).
-We compute average the $\varepsilon$-values of the pair-wise DP guarantees, i.e., the average of the DP-$\varepsilon$’s between all pairs $(i,j)$ over the whole graph, when $\delta=10^{-5}$.
-Interestingly, the bound we obtain using Thm.
-4 for the averaging without secure summation becomes tighter than the bound for the averaging with secure summation (Thm.
-3) as $T$ grows.
-This suggests that the bound of Thm.
-3 can be improved further.
-Notably, both bounds represent clear improvements over the LDP bound.
-Figure 2: DP amplification for adaptive compositions over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=50$ and $p=0.2$.
-Decentralized Learning for Logistic Regression.
-We distribute the MNIST (LeCun,, 1998) dataset IID to 100 nodes and each node is minimizing the cross-entropy loss for the logistic regression model using gradient descent.
-The nodes are connected based on an adjacency matrix $A\in\{0,1\}^{n\times n}$ that corresponds to a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$.
-We compute the average DP-$\varepsilon$ values of all pairs $(i,j)$ (DP of node $j$ from the point of view of node $i$) over the whole graph, when $\delta=10^{-5}$.
-The messages of the nodes are locally DP updated models and nodes carry out gossip averaging over the closed neighborhoods, i.e., the row-stochastic gossip matrix $W=\Lambda^{-1}(A+I_{n})$, where $\Lambda$ is a diagonal matrix with $\Lambda_{ii}=\left|\bar{N_{i}}\right|^{-1}$, $i\in[n]$.
-Figure 3: Decentralized learning for logistic regression on MNIST dataset over a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$.
-The gossip matrix is the row-stochastic matrix corresponding to averaging over the closed neighborhood of each node.
-5 Conclusions and Outlook
-In this work, we have given the first privacy amplification results for DP gossip averaging algorithms on graphs that are able to accurately capture the privacy amplification arising from all the noise injected in the system.
-As a future work, it will be interesting to see, whether the strong results for non-adaptive compositions that use secure summation can be translated to bounds for adaptive compositions.
-From computational perspective, a necessary future task will be to speed up linear algebraic subroutines for evaluating the total sensitivities for large and possibly sparse graphs.
-Also, the proposed framework is likely applicable to other decentralized optimization methods exhibiting a similar linear structure as the gossip averaging considered here.
+6 Conclusions
+We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that secure-summation-based averaging has asymptotic sensitivity comparable to central aggregation.
+Future work includes accelerating the linear-algebra routines needed to compute total sensitivities on large or sparse graphs and analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees.
 Appendix A Proof of Lemma 6
-Lemma 1.
-Let $\sigma>0$ and $A\in\mathbb{R}^{m\times n}$.
-Suppose $\mathcal{M}(D)$ is a projected Gaussian mechanism of the form given in Eq.
-(2.1).
-Let $D$ and $D^{\prime}$ be two datasets such that $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, where $\mathrm{Range(A)}$ denotes the subspace spanned by the columns of $A$.
-Then, for all $\alpha\geq 0$,
-$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{\alpha}\left(\mathcal{N}\big{(}\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2},\sigma^{2}\big{)}\,||\,\mathcal{N}(0,\sigma^{2})\right).$
-i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
 Proof.
@@
 $=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+V_{r}^{\top}Z\,||\,V_{r}^{\top}Z\right)$
-$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
+[…]
 where $\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2}I_{r})$ and where before the third last equality we have carried out multiplication from the left by $U_{r}^{\top}$ and before the second last equality we have carried out multiplication from the left by $\Sigma_{r}^{-1}$.
@@
 We see that for all $\varepsilon\in\mathbb{R}$,
-$H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
+[…]
 where the right-hand side gives the tight $\delta(\varepsilon)$ for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$.
@@
 Appendix B Proof of Lemma 7
-Lemma 1.
-Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality).
-Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
-$\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
-where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
-$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$
 Proof.
-We now need to bound the hockey-stick divergence
-$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right).$
-Recall that from the recursion (3.1) it follows that
-$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix},$
-where
-$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
-$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
-$\vdots$
-$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
-I.e.,
-$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T},$
-where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
-$\widehat{x}_{T}=\begin{bmatrix}x_{0}\\ x_{1}\\ \vdots\\ x_{T-1}\end{bmatrix}.$
-similarly,
-$\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})=H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T},$
-where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
-$\widehat{x}_{T}^{\prime}=\begin{bmatrix}x_{0}^{\prime}\\ x_{1}^{\prime}\\ \vdots\\ x_{T-1}^{\prime}\end{bmatrix}$
-for some $\widehat{x}_{T}^{\prime}\in\mathbb{R}^{T\cdot n}$ such that
-$\widehat{x}_{T}-\widehat{x}_{T}^{\prime}=\begin{bmatrix}e_{j}\Delta_{1}\\ e_{j}\Delta_{2}\\ \vdots\\ e_{j}\Delta_{n}\end{bmatrix},$
-where $e_{j}$ denotes the $j$th standard basis vector in $\mathbb{R}^{n}$ and $\left|\Delta_{i}\right|\leq 1$ for all $i\in[T]$.
 From the translational invariance of the hockey-stick divergence, we have
@@
 $=$ $H_{\alpha}\left(V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+V_{r}^{\top}\widetilde{u}_{T}\,||\,V_{r}^{\top}\widetilde{u}_{T}\right)$
-$=$ $H_{\alpha}\left(V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+\widehat{u}\,||\,\widehat{u}\right)$
-where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$.
-Denoting $\Delta x:=\widehat{x}_{T}-\widehat{x}_{T}^{\prime}$, we see that $-1\leq(\Delta x)_{i}\leq 1$ for all $i\in[T]$.
-For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain $[0,1]^{T}$.
-We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in the convex and compact domain $[0,1]^{T}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$.
-Thus, the optimum is found from somewhere at the boundary, i.e., it holds that $\left|(\Delta x)_{i}\right|=1$ for all $i\in[T]$.
-Moreover, since the rows of $H_{T}$ are vectors with non-negative entries, $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is maximized when $\Delta x=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$.
-∎
-Appendix C Proof of Thm. 3
-Theorem 1.
-Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$.
-Denote
-$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
-Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
+[…]
+[…]
+where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
+Appendix C Proof of Thm. 8
 Proof.
-Suppose $D\simeq_{j}D^{\prime}$.
-Denote $S_{j}=S(\bar{N}_{j})$.
-We need to carry out the analysis for the view $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
-$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$
-$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
-$\vdots$
-$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})+\cdots+W(x_{T-1}+u_{T-1})\big{)},$
-where $[x_{i}]_{k}$ depends on the state $[\theta_{i}]_{k}$ (or could depend also on $[\theta_{i-1}]_{k},[\theta_{i-2}]_{k},\ldots$).
-The only source of randomness in $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ is that of $u_{0},\ldots,u_{T-1}$.
-To carry out the DP analysis, we "roll out" by starting the integration of the hockey-stick divergence from the last noise vector $u_{T-1}$.
-This is similar to the composition analysis using RDP (Mironov,, 2017) and using dominating pairs of distributions (Zhu et al.,, 2022).
-In particular, the proof is similar to the proof of (Thm.
-10, Zhu et al.,, 2022).
-The general case can be illustrated using with the case $T=2$. When $T=2$, we need to analyze the mechanism
-$\mathcal{M}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},$
-where $[x_{1}]_{k}$ depends on the state $[\theta_{1}]_{k}$ for all $k\in[n]$.
-The essential observation here is that when $D\simeq_{j}D^{\prime}$, the the states of the nodes outside of the closed neighborhood $\bar{N}_{j}$ are obtained by applying the same post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\bar{N}_{j})$.
-More specifically, given the view of the nodes inside the closed neighborhood $\bar{N}_{j}$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the nodes outside of $\bar{N}_{j}$, the vectors $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution, since they consist of linear combinations of the entries of $y_{1}$ and of entries of $\theta_{1}$ outside of $\bar{N}_{j}$ which are distributed similarly in case of $D$ and $D^{\prime}$.
-Therefore, when starting to integrate with respect to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and we can also fix all the terms in $Wx_{1}$ that are evaluated using the states of the nodes outside of $\bar{N}_{j}$.
-What remains in $Wx_{1}$ are the terms evaluated using $y_{1}$, and thus we can also carry out an adaptive analysis.
-We then roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
-$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
-where
-$\mathcal{M}_{1}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}$
-and
-$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+Wu_{1}\big{)}\end{bmatrix}.$
-Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
-$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
-where
-$\mathcal{M}_{2}(D)=\begin{bmatrix}S_{j}\big{(}W(e_{j}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(e_{j}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}\big{(}W^{2}+W\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
-and
-$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}Wu_{0}\big{)}\\ S_{j}\big{(}W^{2}u_{0}+Wu_{1}\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
-In case $T=3$, we can use the same reasoning, and similarly, given the observations $y_{1}$ and $y_{2}$, we can freeze the terms $S_{j}\big{(}W^{3}(x_{0}+u_{0})$ and $S_{j}W^{2}(x_{1}+u_{1})$ when integrating w.r.t.
-$u_{2}$.
-The general case also follows from this reasoning.
-Appendix D Proof of Thm. 4
-Theorem 1.
-Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
-$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
-Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
+For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain
+$\mathcal{D}_{j}:=\{\alpha\otimes e_{j}:\alpha\in[-1,1]^{T}\}$
+We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in $\mathcal{D}_{j}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$.
+Thus, the optimum is found from somewhere at the boundary of $\mathcal{D}_{j}$, i.e., it holds that $\Delta x=\alpha\otimes e_{j}$, where $\alpha=\{-1,1\}^{T}$.
+For the inequality, we use the trivial inequality
+$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq\mathbf{1}_{T}^{\top}\left|V_{r}V_{r}^{\top}\right|\mathbf{1}_{T}.$
+Appendix D Proof of Lemma 11
 Proof.
-The proof goes similarly as the proof of Thm. 3 and can be illustrated using the case $T=2$.
-Given the view of the node $j$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the rest of the nodes, the vectors $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution.
-Therefore, when starting to integrate w.r.t.
-to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
-$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
-where
-$\mathcal{M}_{1}(D)=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}$
-and
-$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+u_{1}\big{)}\end{bmatrix}$
-Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
-$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
-where
-$\mathcal{M}_{2}(D)=\begin{bmatrix}e_{j}^{\top}(e_{j}+u_{0})\\ e_{j}^{\top}\big{(}W(e_{j}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}1\\ e_{j}^{\top}\big{(}W+I\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
-and
-$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}Wu_{0}\\ e_{j}^{\top}\big{(}Wu_{0}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
-The general case follows from the same reasoning, similarly as in the proof of Thm. 3.
-Appendix E Results for Time-Varying Graphs
-We see from the proofs that all the results can be straightforwardly generalized to time-varying gossip matrices $W_{t}$, $t\in[T]$.
-For example, in case of the non-adaptive compositions with secure summation, we would have Theorem 1 with
-$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}W_{0}e_{j}\\ e_{i}^{\top}(W_{1}\cdot W_{0}+W_{0})e_{j}\\ \vdots\\ e_{i}^{\top}\left(\prod\limits_{i=0}^{T-1}W_{i}+e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}+\cdots+W_{0}\right)e_{j}\end{bmatrix}$
-and
-$H_{T}:=\begin{bmatrix}e_{i}^{\top}W_{0}&0&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}&e_{i}^{\top}W_{1}&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}\cdot W_{2}&e_{i}^{\top}W_{1}\cdot W_{2}&e_{i}^{\top}W_{2}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}\prod\limits_{i=0}^{T-1}W_{i}&e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}&\dots&e_{i}^{\top}W_{T-2}\cdot W_{T-1}&e_{i}^{\top}W_{T-1}\end{bmatrix}$
-The vector $\widetilde{x}_{T}$ and the matrix $H_{T}$ of Theorems 3 and 4 can be modified analogously.
-Appendix F Details for Experiments
-F.1 Additional Details
-  • For each model, both when using the DP Decentralized Averaging and DP-FedAvg, we optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$.
-  • In case of the gossip averaging, the test accuracies are evaluated on a model of a randomly chosen node.
-  • The experiments on the logistic regression including all the learning rate tuning took around 48 hours using 8 RTX 3080 GPUs.
-F.2 Baseline Method for Experiments
-As a baseline algorithm, we consider the full batch version of the DP Federated Averaging algorithm depicted in Alg. 11. (Ponomareva et al.,, 2023; McMahan et al.,, 2017)
-Algorithm 1 Differentially Private Federated Averaging (DP-FedAvg) with Record-Level DP
-1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise parameter $\sigma$, initial global model $w^{0}$
-2:  for each round $t=0$ to $T-1$ do
-3:     for each client $i$ in parallel do
-4:        Initialize local model: $w_{i}^{t}\leftarrow w^{t}$
-5:        Perform DP gradient descent step on $D_{i}$, with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma/\sqrt{n}$ obtaining updated model $w_{i}^{t}$
-6:        Compute model update: $\Delta_{i}^{t}\leftarrow w_{i}^{t}-w^{t}$
-7:     end for
-8:     Server aggregates updates:
-$\tilde{\Delta}^{t}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}^{t}.$
-9:     Update global model:
-$w^{t+1}\leftarrow w^{t}+\tilde{\Delta}^{t}$
-and distribute it to clients.
-10:  end for
-11:  Output: Final global model $w^{T}$
-F.3 Gossip Averaging Algorithm
-In the experiments, we consider the full decentralized DP gradient descent shown in Alg. 10.
-Algorithm 2 Differentially Private Gossip Averaging (DP-GossipAvg)
-1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise multiplier $\sigma$, gossip matrix $W\in\mathbb{R}^{n\times n}$, initial local models $\{w_{i}^{0}\}_{i=1}^{n}$
-2:  for each round $t=0$ to $T-1$ do
-3:     for each client $i$ in parallel do
-4:        Perform local DP gradient descent on $D_{i}$ with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma$ to obtain updated local model $w_{i}^{t}$
-5:     end for
-6:     for each client $i$ in parallel do
-7:        Gossip averaging:
-$w_{i}^{t+1}\leftarrow\sum_{j=1}^{n}W_{ij}w_{j}^{t}$
-8:     end for
-9:  end for
-10:  Output: Final local models $\{w_{i}^{T}\}_{i=1}^{n}$
+Since $W$ is primitive and row-stochastic, $1$ is a simple eigenvalue with right eigenvector $\mathbf{1}$ and left eigenvector $\pi^{\top}$, and all other eigenvalues satisfy $|\lambda|<1$ (Perron-Frobenius).
+Choose a basis matrix S=[ 1B]S=[\ $\mathbf${1}\ \ B\ ] for some $B\in\mathbb{R}^{n\times n-1}$ so that the first row of $S^{-1}$ is $\pi^{\top}$.
+Then
+$S^{-1}WS\;=\;\begin{bmatrix}1&0\\ 0&T^{\prime}\end{bmatrix}$,
+and taking powers gives
+$W^{t}\;=\;S\begin{bmatrix}1&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\;=\;\underbrace{\mathbf{1}\,\pi^{\top}}_{=\ \Pi}\;+\;\underbrace{S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}}_{=:R^{t}}.$
+To bound $\|R^{t}\|$, fix any matrix norm $\|\cdot\|$.
+Since $\rho(T^{\prime}):=\max\{|\lambda|:\lambda\in\mathrm{spec}(T^{\prime})\}<1$, the Gelfand formula implies that for any $\rho$ with $\rho(T^{\prime})<\rho<1$ there exists $C<\infty$ such that $\|(T^{\prime})^{t}\|\leq C\,\rho^{t}$ for all $t\geq 1$.
+Therefore
+$\|R^{t}\|\;=\;\left\|S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\right\|\;\leq\;\|S\|\,\|S^{-1}\|\,\|(T^{\prime})^{t}\|\;\leq\;\big(\|S\|\,\|S^{-1}\|\;C\big)\,\rho^{t}.$
+Setting $M:=\|S\|\,\|S^{-1}\|\,C$ gives the claim. ∎
+Appendix E Proof Theorem 13
+Proof.
+First, from the compact SVD we see that for any matrix $A$, the product $A^{+}A$ is the orthogonal projector onto $\mathrm{row}(A)$.
+Set $P_{T}:=H_{T}^{+}H_{T}$ to be the orthogonal projection onto row space of $H_{T}$.
+Let $c\in\{\pm 1\}^{T}$ and define $z:=c\otimes e_{j}\in\mathbb{R}^{nT}$.
+Our strategy is to study the sensitivity
+$\Delta_{T}(c,j\to i)=\|P_{T}z\|_{2}$
+and derive a bound for it that is independent of $c$.
+Because $W$ is primitive and row-stochastic, by Lemma 11 there exist constants $M>0$ and $\rho\in(0,1)$ such that
+$W^{t}=\Pi+R^{t},\qquad\Pi:=\mathbf{1}\pi^{\top},\qquad\|R^{t}\|_{2}\leq M\rho^{t}\quad\forall t\geq 0.$ (E.1)
+The $(t,s)$ block of $H_{T}$ (with $0\leq s\leq t\leq T-1$) is given by
+$H_{T}(t,s)=e_{i}^{\top}W^{t-s}=e_{i}^{\top}\Pi+e_{i}^{\top}R^{t-s}=\pi^{\top}+e_{i}^{\top}R^{t-s}.$
+Define the stationary part of $H_{T}$ as $\bar{H}_{T}$ such that blocks of $\bar{H}_{T}$ are given as
+$\bar{H}_{T}(t,s)=\pi^{\top}\quad\text{ for }s\leq t,\qquad\bar{H}_{T}(t,s)=0\quad\text{for }s>t.$
+Furthermore, define the transient part
+$K_{T}:=H_{T}-\bar{H}_{T}.$
+Thus
+$H_{T}=\bar{H}_{T}+K_{T}.$
+We clearly have that
+$\mathrm{row}(\bar{H}_{T})=\{(\alpha_{0}\pi,\alpha_{1}\pi,\dots,\alpha_{T-1}\pi)\in\mathbb{R}^{nT}:\alpha_{t}\in\mathbb{R}\}$
+and we can take as an orthonormal basis for $\mathrm{row}(\bar{H}_{T})$ for example $\{e_{1}\otimes\pi/\|\pi\|_{2},\ldots,e_{T}\otimes\pi/\|\pi\|_{2}\}.$ Therefore the orthogonal projection onto $\mathrm{row}(\bar{H}_{T})$ is given by
+$\bar{P}_{T}z=\sum_{\ell=1}^{T}\langle e_{\ell}\otimes\pi/\|\pi\|_{2},c\otimes e_{j}\rangle=c\;\otimes\;\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi.$
+and we have that
+$\|\bar{P}_{T}z\|_{2}^{2}=\sum_{t=0}^{T-1}\left\|c_{t}\,\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi\right\|_{2}^{2}=T\,\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
+We next study the transient part and show using standard perturbation theory for projection matrices that the distance between the projectors $P_{T}$ and $\bar{P}_{T}$ satisfies
+$\|P_{T}-\bar{P}_{T}\|_{2}=O(1/\sqrt{T}).$
+This rate is derived using the Wedin theorem which states that the distance between the row spaces $\mathrm{row}(H_{T})$ and $\mathrm{row}(\bar{H}_{T})$ is a function of $\|K_{T}\|_{2}$ (since the perturbation $H_{T}-\bar{H}_{T}=K_{T}$) and the smallest non-zero singular value of $\bar{H}_{T}$, $\sigma_{\min}(\bar{H}_{T})$.
+From the Toeplitz structure it follows that clearly $\|K_{T}\|_{2}$ equals the 2-norm of the first column block of $K_{T}$.
+Moreover, from the transient behavior stated in Eq.
+(E.1) and a geometric sum it follows that $\|K_{T}\|_{2}=O(1)$.
+We see that
+$\bar{H}_{T}=U\otimes\pi^{\top}$
+for
+$U=\begin{pmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{pmatrix}$.
+The singular values of a Kronecker product $A\otimes B$ are the products of the singular values of $A$ and $B$. Thus, the non-zero singular values of $\bar{H}_{T}$ are:
+$\sigma_{k}(\bar{H}_{T})=\sigma_{m}(U)\cdot\sigma_{l}(\pi^{\top}).$
+Since $\pi^{\top}$ has rank $1$, $\sigma_{1}(\pi^{\top})=\|\pi^{\top}\|_{2}=\|\pi\|_{2}.$ We also know that the smallest singular value of $U$ is given by
+$\sigma_{\min}(U)=\sqrt{T}.$
+Thus, $\sigma_{\min}(\bar{H}_{T})=\sqrt{T}\|\pi\|_{2}$, and by the Wedin theorem we have that
+$\|P_{T}-\bar{P}_{T}\|_{2}\leq\frac{2\|K_{T}\|_{2}}{\sigma_{\min}(\bar{H}_{T})}=O(1/\sqrt{T}).$
+Let $\mathcal{E}_{T}=P_{T}-\bar{P}_{T}$, so $\|\mathcal{E}_{T}\|_{2}=O(1/\sqrt{T})$.
+We expand $\|P_{T}z\|_{2}^{2}$ using $P_{T}=\bar{P}_{T}+\mathcal{E}_{T}$. Recall $z=c\otimes e_{j}$ and $\|z\|_{2}=\sqrt{T}$:
+$\|P_{T}z\|_{2}^{2}=\|(\bar{P}_{T}+\mathcal{E}_{T})z\|_{2}^{2}=\|\bar{P}_{T}z\|_{2}^{2}+2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)+\|\mathcal{E}_{T}z\|_{2}^{2}.$ (E.2)
+Using a norm bound on $\mathcal{E}_{T}$, we have
+$|2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)|$ $\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}z\|_{2}$
+$\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}\|_{2}\|z\|_{2}$
+$\leq 2\left(\sqrt{T}\frac{\pi_{j}}{\|\pi\|_{2}}\right)\cdot O(1/\sqrt{T})\cdot\sqrt{T}=O(\sqrt{T}).$
+Furthermore,
+$\|\mathcal{E}_{T}z\|_{2}^{2}\leq\|\mathcal{E}_{T}\|_{2}^{2}\|z\|_{2}^{2}=O(1/T)\cdot T=O(1).$
+Substituting these bounds back into Eq. (E.2):
+$\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
+This holds for any choice of $c\in\{\pm 1\}^{T}$. The maximum is thus:
+$(\Delta_{T}(j\to i))^{2}=\max_{c\in\{\pm 1\}^{T}}\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
+Dividing by $T$ and taking the limit:
+$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\lim_{T\to\infty}\left(\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(1/\sqrt{T})\right)=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$

### v3

@@
 Abstract
-Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance.
-However, achieving differential privacy (DP) guarantees in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes.
-We present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph and especially consider the averaging over nearest neighbors with secure summation and individual node-wise views.
-Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes across different scenarios.
-In case the gossip averaging happens via secure summation, we show that the Rényi DP parameter growth is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
+Achieving differential privacy (DP) guarantees in fully decentralized machine learning is challenging due to the absence of a central aggregator and varying trust assumptions among nodes.
+We present a framework for DP analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph.
+We present an analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes.
+Our main contribution is showing that the DP guarantees are those of a Gaussian mechanism, where the growth of the squared sensitivity is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
+As an application of the sensitivity analysis, we show that the excess risk of decentralized private learning for strongly convex losses is asymptotically similar as in centralized private learning.
 1 Introduction
@@
 Fully decentralized gradient-based optimization methods are typically distributed variants of gradient descent and can be broadly classified into two categories.
-In random walk-based methods (rw1; rw2; rw3), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix.
+In random walk-based methods (Lopes and Sayed, 2007; Johansson et al., 2009; Mao et al., 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix.
 The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor.
@@
 However, such sequential update dynamics inherently limits parallelism and scalability.
-In gossip averaging-based methods (gossip2; gossip1), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states.
+In gossip averaging-based methods (Boyd et al., 2006; Dimakis et al., 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states.
 Each node updates its parameters by averaging over the received messages.
@@
 DP in Centralized FL.
-In centralized FL (FLbook), distributed DP (UllahCKO23; google_flblog_2023; FLDP3) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (truex2019hybrid; Erlingsson2019; bell2020secure).
+In centralized FL (Kairouz et al., 2021c), distributed DP (Ullah et al., 2023; Kairouz et al., 2021a) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al., 2019; Erlingsson et al., 2019; Bell et al., 2020).
 This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
 DP in Decentralized Optimization.
-The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (dlattack4; dlattack3; dlattack1; dlattack2).
-Early decentralized DP methods achieve privacy by locally perturbing gradients or models (HMV2015; BGTT18; XZW2022), relying on local DP guarantees.
-This often imposes a poor utility-privacy tradeoff compared to centralized DP (ChanSS12).
-To address this, recent works propose pair-wise network DP (networkdp3; networkdp2; networkdp1), where privacy loss is analyzed between pairs of nodes.
+The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al., 2025; Zhu et al., 2019; Mrini et al., 2024; Pasquini et al., 2023).
+Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al., 2015; Bellet et al., 2018a; Xu et al., 2022), relying on local DP guarantees.
+This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al., 2012).
+To address this, recent works propose pair-wise network DP (Cyffers et al., 2024, 2022; Cyffers and Bellet, 2022), where privacy loss is analyzed between pairs of nodes.
 These relaxations give better utility while accounting for the structure of decentralized communication.
-networkdp2 give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees.
-Recently, bellet2025unified considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (kairouz2021practical; pillutla2025correlated) which leads to considerably more accurate privacy analysis than that of networkdp2.
+Cyffers et al.
+(2022) give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees.
+Recently, Bellet et al.
+(2025) considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025) which leads to considerably more accurate privacy analysis than that of Cyffers et al.
+(2022).
 Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
@@
 This also provides tools to analyze the network DP guarantees in different threat scenarios.
-In addition to considering the threat model of (networkdp2), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
+In addition to including the threat model of (Cyffers et al., 2022), where the neighboring nodes see the plain messages sent by their neighbors, our linear systems view allows analyzing algorithms that incorporate a secure summation protocol between neighboring nodes.
 Secure Aggregation in Decentralized Learning.
-While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored.
-In current gossip-based systems (networkdp2), each node receives messages directly from its neighbors and can, in principle, view their full content.
+While the privacy amplification effect coming from summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored.
+In current gossip-based systems (Cyffers et al., 2022), each node receives messages directly from its neighbors and can, in principle, view their full content.
 As we show, summation over a node’s neighborhood can give a meaningful privacy amplification.
-Several studies have proposed integrating MPC primitives into decentralized learning workflows (jayaraman2018distributed; lian2018asynchronous; jeon2021privacy).
-Concrete protocols for decentralized secure aggregation have been developed in recent works (sabater2022gopa; dlsa1; biswas2024).
-The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (nicolas2024differentially).
+Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al., 2018; Lian et al., 2018; Jeon et al., 2021).
+Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al., 2022; Pereira et al., 2024; Biswas et al., 2024).
+The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al., 2024).
 Our Contributions:
@@
 Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
-  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
-  • We show that in case secure summation is used to carry averaging over the nearest neighbors, the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (networkdp2).
+  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed via the gossip matrix and the linear systems perspective we consider.
+  • We show that in case of a single node view (e.g., that of secure summation based averaging protocols), the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (Cyffers et al., 2022).
+  • As an application of the sensitivity analysis, we combine the privacy bounds with the utility analysis of noisy decentralized learning (Koloskova et al., 2020) and show that the excess risk of strongly convex losses is asymptotically similar to the centralized private learning.
+Relation to Recent Work by (Bellet et al., 2025).
+Recent work by (Bellet et al., 2025) builds on the linear systems formulation of private gossip algorithms and further represents it as a matrix mechanism (Kairouz et al., 2021b; Pillutla et al., 2025).
+Using this formulation, they show that the privacy guarantees extend to adaptive compositions, which we also rely on in this work.
+Our contributions are complementary.
+While (Bellet et al., 2025) focus on adaptive privacy and general sensitivity bounds via the matrix mechanism view, we work directly in the linear systems formulation to derive analytic characterizations of both privacy and convergence for decentralized learning, which require this representation and are not obtained in their analysis.
+Our results show that private decentralized learning can match centralized methods in both privacy and utility.
 2 Background
@@
 We say that a mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable.
-Definition 1 (dwork_et_al_2006).
+Definition 1 (Dwork et al. 2006).
 Let $\varepsilon\geq 0$ and $\delta\in[0,1]$.
@@
 The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
-Lemma 2 (balle2018subsampling).
+Lemma 2 (Balle et al. 2018).
 A mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if and only if, $\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\delta$.
@@
 $\mathcal{M}^{(T)}(D)=\big(\mathcal{M}_{1}(D),\mathcal{M}_{2}(\mathcal{M}_{1}(D),D),\ldots,\mathcal{M}_{T}(\mathcal{M}_{1}(D),\ldots,\mathcal{M}_{T-1}(D),D)\big).$
-The results of (balle2018gauss) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism.
-Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by dong2022gaussian.
-Lemma 4 (dong2022gaussian).
+The results of (Balle and Wang, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism.
+Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al.
+(2022).
+Lemma 4 (Dong et al. 2022).
 Consider an adaptive composition of $T$ Gaussian mechanisms, each with $L_{2}$-sensitivity $\Delta$ and noise scale parameter $\sigma$. The adaptive composition is $(\varepsilon,\delta)$-DP for
 $\delta(\varepsilon)=\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}+\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right)-e^{\varepsilon}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}-\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right).$
+An alternative and often more convenient way to analyze privacy guarantees is via Rényi Differential Privacy (RDP) (Mironov, 2017), which is based on Rényi divergence.
+For $\alpha>1$, the Rényi divergence of order $\alpha$ between two distributions $P$ and $Q$ is defined as
+$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\!\left[\left(\frac{P}{Q}\right)^{\alpha}\right].$
+A mechanism $\mathcal{M}$ is said to satisfy $(\alpha,\varepsilon)$-RDP if for all neighboring datasets $D\sim D^{\prime}$, $D_{\alpha}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq\varepsilon.$ The Gaussian mechanism with $\ell_{2}$-sensitivity $\Delta$ and noise variance $\sigma^{2}$ satisfies $(\alpha,\varepsilon)$-RDP with $\varepsilon=\frac{\alpha\,\Delta^{2}}{2\sigma^{2}}.$
+We next give a result for a certain projected Gaussian mechanism, where the noise vector is projected to a smaller-dimensional subspace.
 2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
-In our results, the network DP guarantees become those of a projected Gaussian mechanism of the form
+In our results, the network DP guarantees become those of a Gaussian mechanism of the form
 $\mathcal{M}(D)=f(D)+AZ,$ (2.1)
@@
 To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$.
-We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (golub2013matrix).
+We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan, 2013).
 Definition 5 (Compact SVD and Moore–Penrose Pseudoinverse).
@@
 The Moore–Penrose pseudoinverse of $A$, denoted $A^{+}$, is defined via compact SVD as
-$A^{+}=V_{r}\Sigma_{r}^{+}U_{r}^{\top}$
+$A^{+}=V_{r}\Sigma_{r}^{+}U_{r}^{\top},$
 where $\Sigma_{r}^{+}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with entries $\frac{1}{\sigma_{1}},\dots,\frac{1}{\sigma_{r}}$, i.e., $\Sigma_{r}^{+}=\mathrm{diag}\left(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}},\dots,\frac{1}{\sigma_{r}}\right).$
@@
 2.3 Decentralized Learning and Network Differential Privacy
-We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an adjacency matrix
-$A\in\{0,1\}^{n\times n},\qquad A_{ij}=1\;\text{iff}\;(i,j)\in E.$
-We denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
+We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes.
+The graph structure is encoded by an $n\times n$ boolean adjacency matrix and we denote the nodes by indices $n]=\{1,\dots,n\}$.
+For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
 $N_{i}=\{j\in[n]:(i,j)\in E\},\qquad\bar{N}_{i}=N_{i}\cup\{i\}.$
-Local Data and Learning Objective.
-The global dataset is partitioned across nodes:
+Local Data and Learning Objective. The global dataset is partitioned across nodes:
 $D=\bigcup_{i=1}^{n}D_{i},$
@@
 At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$.
-For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by bellet2025unified as described in Section 3.4.
+For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by Bellet et al.
+(2025) as described in Section A.
 Dataset Adjacency.
-We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
+We adopt a record-level adjacency relation.
+Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
 $D\sim_{i}D^{\prime}\quad\iff\quad D_{j}=D^{\prime}_{j}\ \text{for all }j\neq i.$
-Mechanism and Observations.
-A decentralized mechanism is a randomized mapping
-$\mathcal{M}:D\mapsto\text{(all messages exchanged during the protocol)}.$
-For a subset of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{A}$ is denoted
+Mechanism and View. We define a decentralized mechanism as a randomized mapping
+$\mathcal{M}:D\mapsto\text{(all variables exchanged during the protocol)}.$
+For a set of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{M}$ from the point-of-view of $\mathcal{A}$, i.e., the part of $\mathcal{M}(D)$ seen by nodes in $\mathcal{A}$, is denoted by
 $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A}).$
@@
 where $e_{k}$ is the $k$th standard basis vector.
-Network Differential Privacy.
-Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
+Network Differential Privacy. Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
 $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\;\simeq_{(\varepsilon,\delta)}\;\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$
 We assume honest-but-curious nodes: they follow the protocol but analyze observed messages.
-As in (networkdp2), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
-Gossip-Based Averaging.
-Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
+As in (Cyffers et al., 2022), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
+Gossip-Based Averaging. Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
 $\theta_{t+1}=Wx_{t},$
 where $W\in\mathbb{R}^{n\times n}$ is the gossip matrix. We call $W$:
-$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},\qquad\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
-For the error analysis we also consider the definition of primitive gossip matrices.
-Definition 2.1 (Primitive Matrix).
+$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},$
+and
+$\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
+Primitive gossip matrices have vanishing transient components, and we focus on this class in our asymptotic analysis.
+Definition 7 (Primitive Matrix).
 A matrix $W\in\mathbb{R}^{n\times n}$ is called primitive if there exists $k\in\mathbb{N}$ such that
-$W^{k}>0\quad\text{(entrywise)}.$
+$W^{k}>0\quad\text{(entry-wise)}.$
 Equivalently, a row-stochastic matrix is primitive if and only if the induced Markov chain is irreducible and aperiodic.
@@
 $W^{t}\to\mathbf{1}\pi^{\top}\qquad\text{as }t\to\infty.$
+Remark 8.
+A primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value strictly less than 1.
+We repeatedly use the following auxiliary lemma to bound the transient part of $W$.
+Lemma 9.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Then there exists $\rho\in(0,1)$ and a symmetric matrix $R$ such that
+$W=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}+R,$
+$\mathbf{1}_{n}^{\top}R=0$, and for all integers $t\geq 0$, $\|R^{t}\|_{2}\leq\rho^{t}.$
 3 Privacy Analysis
 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
-We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., antoulas2005approximation).
-In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
+We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas, 2005).
+In particular, we consider the systems without the feedthrough term in which case their dynamics are described by the equations
 $\theta_{t+1}$ $=A_{t}\theta_{t}+B_{t}u_{t},$
@@
 where at each time step $t$, $\theta_{t}$ represents the state vector, the matrix $A_{t}$ the state transition matrix, $B_{t}$ the input matrix which describes how the control input $u_{t}$ at time step $t$ affects the state, and the observation $y_{t}$ is related to the state vector $\theta_{t}$ by the observation matrix $C$.
-We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch.
-4 antoulas2005approximation).
+We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{T}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[T]$ (see, e.g., Ch.
+4 Antoulas, 2005).
 In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
-3.2 Analysis of Non-Adaptive Compositions
-We first show how to analyze DP gossip averaging with secure summation protocols for non-adaptive node-wise functions of data.
+3.2 DP Analysis of Compositions
+We show how to analyze DP gossip averaging algorithms using the linear systems view.
+We first focus on non-adaptive compositions, where the node-wise contributions do not depend on the state variables.
+The connection between the linear systems view and matrix mechanisms as established by Bellet et al.
+(2025) then allows extending all the results to the adaptive case.
 To this end, without loss of generality, consider the task of globally averaging node-wise streams of data.
@@
 Lemma 10.
-Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$.
-Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
+Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $c_{t}\in\mathbb{R}$.
+Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of the node set $\mathcal{A}\subset[n]$, where $(\varepsilon,\delta)$ is given by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
 $\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
 where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
-$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$ (3.2)
-Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 7 directly gives the following result.
+$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)$
+and
+$H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$, (3.2)
+and $S$ is the selector matrix for $\mathcal{A}$.
+Remark 11.
+For example, in case we consider the threat model where all the messages sent by the neighbors are visible to a given node $i$, the selector matrix $S$ can be taken as $S=S(\bar{N}_{i})$ corresponding to the view of a closed neighborhood of node $i$.
+In case of secure summation over the nearest neighbors, we may take $S=e_{i}^{\top}$, i.e., the view of a single node.
+Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 10 gives the following general $(\varepsilon,\delta)$-DP bound the noisy gossip averaging.
 Theorem 12.
 Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$.
+Suppose also that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$.
 Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq.
+(3.2) for some $\mathcal{A}\subset[n]$.
+Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$.
+Let $\widetilde{x}_{T}$ be of the form given in Eq.
 (3.2).
-Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$.
-Let $\widetilde{x}_{T}$ be of the form (3.2) and denote $c=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{T}\end{bmatrix}^{\top}\in\mathbb{R}^{T}$.
 Denote
-$\Delta^{T}_{j\rightarrow i}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
+$\Delta^{T}_{j\rightarrow\mathcal{A}}$ $=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}$
+$=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}$
+$\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
 where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector.
-Then, from the point-of-view of a node $i$, the node $j$ is tightly $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
+Then, from the point-of-view of the node set $\mathcal{A}$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
+As we outline in Appendix A, due to the connection between the linear systems view and the so called matrix mechanisms, as shown by Bellet et al.
+(2025), Theorem 12 holds also in case adaptive cade, i.e., when the node-wise contributions $x_{t}$ depend on the state variable $\theta_{t}$.
+Moreover, from the spherical symmetry of the Gaussian noise it follows that the results generalize to arbitrary dimensions, when the condition $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$, is replaced by 2-norm bound for the sensitivity of node $j$’s contribution.
+We also remark that the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ leads to a tight $(\varepsilon,\delta)$-DP bound.
 3.3 Accounting for Nodes’ Knowledge of Injected Noise
-We have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes.
-In practice, this likely has a small effect on the the DP guarantees.
-However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes.
+We have so far assumed that the noise injected by the nodes belonging to the observing set $\mathcal{A}\in[n]$ contributes to the DP guarantees of the other nodes.
+We can obtain rigorous guarantees also in case they do not contribute, by removing the noise terms corresponding to the observing nodes.
 This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees.
-Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by allouahprivacy is scheme called SecLDP.
+Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by Allouah et al.
+(2024) is scheme called SecLDP.
 For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm.
-8, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq.
+12, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq.
 (3.2) with the $i$th column vector of each $T\times n$-block column removed.
-Similar correction can be carried out also in all the other cases.
-We do this correction in all our experiments.
-3.4 Adaptive Compositions and Connection to Matrix-Mechanism Accounting
-The recent work by bellet2025unified shows that the linear systems perspective given in Theorem 8 can be equivalently analyzed via a so-called matrix mechanisms (kairouz2021practical; pillutla2025correlated).
+For a general adversarial set $\mathcal{A}$, we obtain a complete analogy of Lemma 10 and Thm.
+12 with the matrix blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$.
+3.4 Compute Efficient Upper Bound
+We obtain a compute efficient upper bound for the sensitivity from a quadratic-form representation.
+Since $H_{T}$ has full row rank, $K_{T}:=H_{T}H_{T}^{\top}$ is invertible and for all $v\in\mathbb{R}^{m\cdot T\times 1}$
+$\left\|H_{T}^{+}H_{T}v\right\|_{2}^{2}=(H_{T}v)^{\top}K_{T}^{-1}(H_{T}v).$
+In particular, for $c\in\{-1,1\}^{T}$ and $v(c)=c\otimes e_{j}$, letting $b(c):=H_{T}v(c)$ we have
+$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$
+Moreover, we can construct a matrix $G_{T}\in\mathbb{R}^{mT\times T}$ such that $b(c)=G_{T}c.$ Hence, we can construct an $Tm\times Tm$ matrix $M_{T}:=G_{T}^{\top}K_{T}^{-1}G_{T}\succeq 0$, with which we get the efficiently computable upper bound
+$\Delta_{T}(j\to\mathcal{A})^{2}$ $=\max_{c\in\{\pm 1\}^{T}}c^{\top}M_{T}c$
+$\leq\max_{\|x\|_{2}^{2}=T}x^{\top}M_{T}x$
+$=T\,\lambda_{\max}(M_{T}).$
+We illustrate this bound numerically in Appendix F.
+4 Asymptotic Sensitivity Growth Rate
+We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
+Theorem 13.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic and primitive.
+Let $H_{T}$ be the system matrix given in Lemma 10 with $\mathcal{A}=\{i\}$, i.e., $S=e_{i}^{\top}$ for some $i\in[T]$ (single-node view) and define
+$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
+as in Theorem 12. Then
+$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{1}{n}.$
+Remark 14.
+Notice that $(\Delta_{T}(j\to i))^{2}\sim T/n$ matches the sensitivity of the centralized case when each of the $n$ devices adds independent Gaussian noise with variance $\sigma^{2}$: their average has the output noise variance $n\sigma^{2}$ and as a result the DP analysis is equivalent to that of a Gaussian mechanism with noise scale $\sigma$ and squared sensitivity $T/n$ over a $T$-wise composition.
+Using the proof technique of Thm.
+13, we can also obtain the following sensitivity bound for the case the observing set $\mathcal{A}$ is of size $m$, $m>1$.
+Notice that Theorem 15 does not give a lower bound: as the proof indicates, the tight sensitivity becomes dependent on the gossip matrix $W$ and the observing set $\mathcal{A}$.
+Although the proof of Theorems 15 and 16 are analogous to the proof of Theorem 13 we state all these results separately for readability.
+Theorem 15.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly-stochastic, and primitive.
+Fix a set of observing nodes $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$, and let $S_{\mathcal{A}}\in\mathbb{R}^{m\times n}$ be the row-selector matrix that extracts the coordinates indexed by $\mathcal{A}$ (i.e., $S_{\mathcal{A}}x=x_{\mathcal{A}}$).
+Let $H_{T}$ be the system matrix of Lemma 10 (Eq.
+(3.2)) constructed with observation matrix $S=S_{\mathcal{A}}$.
+For any $j\in\{1,\dots,n\}$ define the sensitivity
+$\Delta_{T}(j\to\mathcal{A}):=\max_{c\in\{\pm 1\}^{T}}\Bigl\|H_{T}^{+}H_{T}\,(c\otimes e_{j})\Bigr\|_{2}.$
+Then, denoting $\rho:=\max\{\left|\lambda_{2}(W)\right|,\left|\lambda_{n}(W)\right|\}$, we have that for all $T\geq 1$,
+$\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T+\frac{2}{(1-\rho)^{2}}\sqrt{\frac{m}{n}\,T}+\frac{2}{(1-\rho)^{4}}$
+which further implies $\lim_{T\to\infty}\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T$.
+Using the proof technique of Thm.
+13, we also get an asymptotic upper bound for the case there are $m$ observing nodes and the noise injected by them does not contribute to the DP guarantees.
+We remark that the proof of Thm.
+16 gives a similar explicit upper bound as in Thm.
+15.
+Theorem 16.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive.
+Fix an observing set $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$.
+Suppose $j\notin\mathcal{A}$ and suppose the noise terms injected by the nodes in $\mathcal{A}$ do not contribute to the DP guarantees.
+Denote by $\bar{\Delta}_{T}(j\to\mathcal{A})$ the sensitivity of Theorem 12 with the blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$.
+Then,
+$\limsup_{T\to\infty}\frac{\bar{\Delta}_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n-m}.$
+4.1 Experimental Illustration of the DP Bounds for Single Node View
+We next consider a numerical illustration of the bound given by Theorem 13 in case a single node is observed.
+First, consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, i.e., each of the $n$ users is connected to each other with probability $p$.
+Second, consider a randomly drawn graph generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes.
+Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees.
+This model produces a structure characterized by heavy-tailed degree distributions.
+In both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
+As Figure 1 illustrates, in case only a single node $i$ is viewed, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ grows asymptotically linearly w.r.t.
+$T$ with the rate $1/n$ as predicted by Thm.
+13.
+Here both $i$ and $j$ are chosen randomly.
+Figure 1: The scaled sensitivity $(\Delta_{T}(j\to i))^{2}/T$ as a function of $T$ for a) randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes.
+We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
+5 Utility Bound for Strongly Convex Losses
+We next show how our sensitivity analysis can be plugged into standard convergence analysis of decentralized optimization (Koloskova et al., 2020) to obtain explicit privacy-utility bounds for private decentralized learning based on gossip averaging.
+For simplicity, we focus on the case of single-node view, however the results can be directly generalized to multi-node view using Theorems 15 and 16.
+Setup. We consider decentralized minimization of the loss over $n$ nodes,
+$f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$
+over a connected undirected graph with a fixed symmetric, doubly-stochastic gossip matrix $W$.
+Assume each $f_{i}$ is $L$-smooth and $\mu$-strongly convex.
+Each node performs full-batch DP gradient descent, followed by gossip averaging:
+$\theta_{i}^{t+\frac{1}{2}}$ $=\theta_{i}^{t}-\eta\,(\nabla f_{i}(\theta_{i}^{t})+\xi_{i}^{t}),$ (5.1)
+$\theta_{i}^{t+1}$ $=\sum_{j=1}^{n}w_{ij}\,\theta_{j}^{t+\frac{1}{2}},$
+where $\xi_{i}^{t}\sim\mathcal{N}(0,\sigma_{\mathrm{DP}}^{2}I_{d})$ are independent across $i,t$.
+Following (Koloskova et al., 2020), we measure utility in terms of the average iterate $\bar{\theta}^{t}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}^{t}.$
+Topology parameter (absolute spectral gap). Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be eigenvalues of $W$ and define the absolute spectral gap
+$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$ (5.2)
+In the decentralized SGD analysis by Koloskova et al. (2020), the effect of the network enters through a consensus-rate parameter $p$; for fixed symmetric $W$ one may take (see Lemma J.1 in Appendix)
+$p=1-\gamma^{2}\;\;\;\text{and hence}\;\;\;\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right).$ (5.3)
+Using Sensitivity Bound to Set DP Noise.
+Under the threat model of a single-node view $S=e_{i}^{\top}$, our Theorem 13 shows that the total sensitivity from node $j$ to observer $i$, $\Delta_{T}(j\to i))^{2}$ is asymptotically $T/n$ for doubly-stochastic $W$.
+We get an explicit upper bound from Thm.
+15:
+$(\Delta_{T}(j\to i))^{2}\;\leq\;\frac{T}{n}+\frac{2}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{2}{\gamma^{4}}.$ (5.4)
+To guarantee $(\varepsilon,\delta)$-DP, we set the Gaussian noise as a function of $\varepsilon$ and $\delta$ via the standard condition (Dwork and Roth, 2014)
+$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\,\sqrt{2\log(1.25/\delta)}}{\varepsilon}.$ (5.5)
+Combining (5.5) with (5.4) gives the per-step DP noise scale
+$\sigma_{\mathrm{DP}}^{2}\;\lesssim\;\frac{\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right).$ (5.6)
+Gradient Noise for Convergence Bounds of (Koloskova et al., 2020).
+Because we use full-batch gradients, the only stochasticity is the injected DP noise, and the gradient-noise variances needed for the results of (Koloskova et al., 2020) are given by
+$\sigma_{i}^{2}:=\mathbb{E}\|\xi_{i}^{t}\|_{2}^{2}=d\,\sigma_{\mathrm{DP}}^{2},$ (5.7)
+which implies the mean noise parameter $\bar{\sigma}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}=d\,\sigma_{\mathrm{DP}}^{2}.$ Following Koloskova et al.
+(2020), we also define the heterogeneity measure
+$\bar{\zeta}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\theta^{\star})\|_{2}^{2}.$
+5.1 Utility Bound for DP Decentralized Learning
+We can now directly apply the analysis by Koloskova et al.
+(2020) by using (5.6) and (5.7).
+We state the result in a form that the topology appears via $\gamma$ and privacy via $(\varepsilon,\delta)$.
+More details are given Appendix J.
+Theorem 17 (Utility of Full-Batch DP-GD Training with Gossip Averaging).
+Assume $f_{i}$ are $L$-smooth and $\mu$-strongly convex, and $W$ is symmetric, doubly-stochastic with absolute spectral gap $\gamma$.
+Run (5.1) for $T$ rounds with constant stepsize $\eta\simeq 1/L$ and Gaussian noise calibrated so that any node $j$ is $(\varepsilon,\delta)$-DP from the point of view of any node $i$, using (5.5) with $\Delta_{T}(j\to i)$ and the finite bound (5.4).
+Then the time-averaged excess risk of the average iterate satisfies
+$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{\theta}^{t})-f^{\star}\right]\;\leq$ $\widetilde{O}\!\bigg(\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n^{2}\,\mu}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{2}\sqrt{T}}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{4}T}$
+$\quad\;+\;\frac{L\,\bar{\zeta}^{2}}{\mu^{2}\,\gamma^{2}\,T^{2}}\;+\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{\gamma}\exp\!\Big(-c\,\gamma\frac{\mu}{L}T\Big)\bigg),$
+[…]
+Proof on a High Level.
+Theorem 17 is obtained by: (i) viewing full-batch DP gradient descent with gossip averaging as decentralized SGD with gradient noise variance $\bar{\sigma}^{2}=d\sigma_{\mathrm{DP}}^{2}$, (ii) applying the strongly-convex convergence bound by Koloskova et al.
+(2020) with network parameter $p$ and substituting $p=\Theta(\gamma)$, and (iii) determining $\sigma_{\mathrm{DP}}$ via our single-node-view sensitivity given in Section 4.
+Interpretation and Comparison to Centralized DP.
+The leading term in (17) is a privacy-induced error term $\Theta\!\left(\frac{d\log(1/\delta)}{\varepsilon^{2}n^{2}\mu}\right),$ which matches the scaling one has in the centralized setting (Bassily et al., 2014, 2019, 2020) (see also Sec.
+4.2, Ponomareva et al., 2023).
+The remaining terms are transient and topology-dependent via the absolute spectral gap $\gamma$, vanishing as $T$ grows.
+Thus, the topology affects how quickly one reaches the privacy-utility baseline of the centralized setting.
+6 Conclusions
+We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that private gossip averaging gives asymptotically privacy guarantees comparable to central aggregation in case of a single-node view which further leads to utility results comparable to those of centralized private convex learning.
+We also analyzed the case of multiple-node view and SecLDP, the scenario where the noise injected by the viewing nodes does not contribute to the privacy guarantees.
+Future work includes analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees and privacy-utility trade-offs.
+Appendix A Adaptive Compositions and Connection to Matrix-Mechanism Accounting
+The recent work by Bellet et al.
+(2025) shows that the linear systems perspective given in Theorem 12 can be equivalently analyzed via a so-called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025).
 This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis.
 We next shortly discuss this connection.
-At a high level, both frameworks reduce an observer’s view to a linear image of a data-dependent signal plus correlated Gaussian noise.
 Different Threat Scenarios.
-The matrix mechanism formulation by bellet2025unified allows analyzing private decentralized optimization methods that have a linear structure.
+The matrix mechanism formulation by Bellet et al.
+(2025) allows analyzing private decentralized optimization methods that have a linear structure.
 Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
-  • The LDP approach where the adversary sees all messages (bellet2018personalized). This corresponds to the setting of Theorem 8 with $S=I_{n}$.
-  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (networkdp2).
+  • The LDP approach where the adversary sees all messages (Bellet et al., 2018b). This corresponds to the setting of Theorem 12 with $S=I_{n}$.
+  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (Cyffers et al., 2022).
 This corresponds to Theorem 12 with a selector matrix of a finite collection of nodes.
-  • SecLDP (allouahprivacy), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees.
-This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Thmeorem 8, see Section 3.3.
+  • SecLDP (Allouah et al., 2024), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees.
+This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Theorem 12, see Section 3.3.
 Privacy Analysis of Matrix Mechanisms.
-The matrix-mechanism analysis by bellet2025unified provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings.
+The matrix-mechanism analysis by Bellet et al.
+(2025) provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings.
 The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
@@
 The privacy analyses of different private optimization methods under varying threat scenarios are then analyzed via the following results.
-Theorem 9 (Bellet et al., 2025, Thm. 5).
+Theorem A.1 (Bellet et al., 2025, Thm. 5).
 For each of the three trust models, LDP, PNDP and SecLDP, and for every linear decentralized learning algorithm, there exist matrices $A$, $B$ and $C$ such that the adversarial view can be written as
-$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (3.3)
+$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (A.1)
 with $A=BC$.
 To quantify privacy, one needs to bound how much $G$ may vary under the adjacency relation while remaining observable through this linear mapping.
-Definition 3.1 (Sensitivity under participation).
+Definition A.2 (Sensitivity under participation).
 Let $B$ and $C$ be two matrices, and let $\Pi$ be a participation scheme for $G$. The sensitivity of $C$ with respect to $B$ under $\Pi$ is defined as
@@
 In other words, $\Pi$ encodes which entries of $G$ correspond to contributions of the same individual across time, ensuring that sensitivity is computed only over data differences permitted by the privacy definition.
-Theorem 10 (Bellet et al., 2025, Thm. 7).
+Theorem A.3 (Bellet et al., 2025, Thm. 7).
 Let $\mathcal{O}_{\mathcal{A}}=AG+BZ$ be the attacker knowledge of a trust model on a linear decentralized learning algorithm, and denote by $\mathcal{M}(G)$ the corresponding mechanism.
@@
 Adaptive Analysis.
-Notably, bellet2025unified extend the existing adaptive analysis of matrix mechanisms (denisov2022improved) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns.
+Notably, Bellet et al.
+(2025) extend the existing adaptive analysis of matrix mechanisms (Denisov et al., 2022) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns.
 The crucial implication is that once the representation
@@
 We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step.
-As shown in (Appendix A.4 bellet2025unified), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm.
-8.
-Equivalence of Sensitivity Bounds.
-We also see from Thm.
-10 that the sensitivity upper bound equals the upper bound given in our Thm.
+As shown in (Appendix A.4 Bellet et al., 2025), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm.
+12.
+Equivalence of Sensitivity Bounds of Thm.
+12 and of (Bellet et al., 2025).We also see from Thm.
+A.3 that the sensitivity upper bound equals the upper bound given in our Thm.
 12 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
 $\left|C^{\top}B^{+}BC\right|=\left|VV^{\top}\right|$
-and the summation in the upper bound of Thm. 10 equals the upper bound given in our Thm. 8 in case the differing node participates in every round.
-4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
-We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
-We first need the following auxiliary lemma.
-Lemma 11.
-Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic. Then there exist constants $M<\infty$ and $\rho\in(0,1)$ such that for all $t\geq 1$
-$W^{t}\;=\;\Pi\;+\;R^{t},\qquad\Pi:=\mathbf{1}\,\pi^{\top},\qquad\|R^{t}\|\;\leq\;M\,\rho^{t},$
-where $\pi^{\top}W=\pi^{\top}$, $\pi>0$, $\sum_{i}\pi_{i}=1$.
-Remark 12.
-A row-stochastic primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value < 1, and as the proof of Lemma 11 shows, $1-\rho$ equals the spectral gap $\lambda_{1}-\max_{i\geq 2}\left|\lambda_{i}\right|$.
-Theorem 13.
-Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic with stationary distribution $\pi>0$, $\sum_{k=1}^{n}\pi_{k}=1$, $\pi^{\top}W=\pi^{\top}$.
-Let $H_{T}$ be the system matrix given in Lemma 7 with $S=e_{i}^{\top}$ (view of a single node, i.e., that of secure summation protocol) and define
-$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
-as in Theorem 8. Then
-$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
-Remark 14.
-Notice that if $W$ is doubly stochastic, $\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}=\frac{1}{n}$ and we asymptotically have the same sensitivity as in case of a central aggregator.
-5 Experiments
-Illustration of the DP Bounds for Secure Summation Protocols. We first consider numerical evaluation of the DP bounds we obtain from Thm. 8.
-First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$.
-We take the gossip matrix $W$ to be a row-stochastic matrix obtained from row-wise averaging.
-Second, we consider a network generated using a preferential-attachment model (barabasi1999emergence) with $n=200$ nodes, initialized from a fully connected core of 5 nodes.
-Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees.
-The gossip matrix $W$ is obtained from row-wise averaging.
-As Figure 1 illustrates, in case of the secure summation, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ for a randomly chosen $j$ grows asymptotically linearly w.r.t. $T$ as predicted by Thm. 13.
-Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: gossip matrix of a preferential-attachment model with $n=200$ nodes, and the scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$.
-We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
-6 Conclusions
-We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that secure-summation-based averaging has asymptotic sensitivity comparable to central aggregation.
-Future work includes accelerating the linear-algebra routines needed to compute total sensitivities on large or sparse graphs and analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees.
-Appendix A Proof of Lemma 6
+and the summation in the upper bound of Thm. A.3 equals the upper bound given in our Thm. 12 in case the differing node participates in every round.
+Appendix B Proof of Lemma 6
 Proof.
@@
 Since $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, it holds that
-$f(D)-f(D^{\prime})=U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big).$ (A.1)
+$f(D)-f(D^{\prime})=U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big).$ (B.1)
 By using the compact SVD of $A$, we have that for all $\alpha\geq 0$:
@@
 […]
-where the right-hand side gives the tight $\delta(\varepsilon)$ for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$.
+where the right-hand side gives the $\delta(\varepsilon)$-privacy profile for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$.
 Further, since $V_{r}$ has orthonormal columns,
 $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ $=\left\lVert V_{r}\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
-$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
+$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2},$
 and the claim follows. ∎
-Appendix B Proof of Lemma 7
+Appendix C Proof of Lemma 9
+Proof.
+Since $W$ is symmetric, $W=Q\Lambda Q^{\top}$ with $Q$ orthogonal and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ real.
+Doubly stochastic implies $\lambda_{1}=1$ with eigenvector $\mathbf{1}$.
+Primitivity implies $|\lambda_{k}|<1$ for all $k\geq 2$.
+Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$ and set R:=WPR:=W-P.
+Then $RP=PR=0$ and $\|R^{t}\|_{2}=\max_{k\geq 2}|\lambda_{k}|^{t}$.
+With $\rho:=\max_{k\geq 2}|\lambda_{k}|\in(0,1)$ the claim follows.
+∎
+Appendix D Proof of Lemma 10
 Proof.
@@
 $H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right)$ $=H_{\alpha}\left(H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T}\right)$
-$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right).$
+$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
+which gives the claim.
+Appendix E Proof of Thm. 12
+Proof.
+The claim is obtained by applying Lemma 6 (DP guarantees of a projected Gaussian mechanism via Moore–Penrose pseudoinverse) to the characterization given by Lemma 10.
+For completeness, we apply the pseudoinverse $H_{T}^{+}$ step by step.
+By Lemma 10, the privacy guarantees are obtained via the hockey-stick divergence $H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$, $\alpha\geq 0$.
 By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
-$H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
+$H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
 $=H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
-$=$ $H_{\alpha}\left(\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
+$=H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
 $=H_{\alpha}\left(V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+V_{r}^{\top}\widetilde{u}_{T}\,||\,V_{r}^{\top}\widetilde{u}_{T}\right)$
@@
 […]
-where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
-Appendix C Proof of Thm. 8
-Proof.
+where
+$c=\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}$
+and $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
 For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain
@@
 For the inequality, we use the trivial inequality
-$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq\mathbf{1}_{T}^{\top}\left|V_{r}V_{r}^{\top}\right|\mathbf{1}_{T}.$
-Appendix D Proof of Lemma 11
+$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|V_{r}V_{r}^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
+where $\left|\cdot\right|$ denotes the element-wise abolute value of a matrix. ∎
+Appendix F Numerical Illustration of the Spectral Estimate of Sec. 3.4
+We next consider a numerical illustration of the upper bound for the sensitivity
+$\Delta^{T}_{j\rightarrow\mathcal{A}}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2},$
+where $H_{T}$ is the system matrix given in Lemma 10.
+We compare the upper bound based on the spectral estimate $\lambda_{\max}(M_{T})$ to a lower bound for the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ we obtain with the choice $c=\mathbf{1}=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$.
+We compare the bounds in the same two cases considered in Section 4.1: we first consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, and second, a network generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes, where each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees.
+And in both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
+As illustrated by Figure 2 the upper bound is in both cases less than 10 percent higher than the actual tight upper bound.
+Figure 2: The upper bound for the squared and scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$ for a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes.
+Appendix G Proof of Theorem 13
+We next give proof of Theorem 13. In the proof, Lemma G.2 (Section G.2) contains a crucial step: it shows that the transient component of the sensitivity can be bounded independent of $T$.
+G.1 Proof of Theorem 13
+Theorem G.1.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Let $S=e_{i}^{\top}$ be the single-node observation matrix. Then for any $j\in\{1,\dots,n\}$,
+$\lim_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}=\frac{1}{n},$
+where
+$\Delta_{T}(j\to i)^{2}:=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}^{2}.$
 Proof.
-Since $W$ is primitive and row-stochastic, $1$ is a simple eigenvalue with right eigenvector $\mathbf{1}$ and left eigenvector $\pi^{\top}$, and all other eigenvalues satisfy $|\lambda|<1$ (Perron-Frobenius).
-Choose a basis matrix S=[ 1B]S=[\ $\mathbf${1}\ \ B\ ] for some $B\in\mathbb{R}^{n\times n-1}$ so that the first row of $S^{-1}$ is $\pi^{\top}$.
+We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
+$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{T},$
+and set
+$K_{T}:=H_{T}H_{T}^{\top}.$
+Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
+$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
+hence
+$\Delta_{T}(j\to i)^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (G.1)
+Denoting $a_{m}:=(W^{m})_{ij}$ for $m\in\mathbb{N}$, we have for $t=1,\dots,T$,
+$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}c_{k}.$ (G.2)
+We start by decomposing $K_{T}$. By Lemma 9, we can write
+$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (G.3)
+and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|$. Using (G.3) together with the block-Toeplitz structure of $H_{T}$, we obtain the decomposition
+$K_{T}=\frac{1}{n}J+G_{T},$ (G.4)
+where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$, with $L$ the $T\times T$ lower-triangular all-ones matrix:
+$L=\begin{bmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
+Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
+$K_{T}\succeq\frac{1}{n}J$
+which further implies that
+$K_{T}^{-1}\preceq nJ^{-1}.$ (G.5)
+Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
+$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
+i.e.,
+$D=\begin{bmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&-1&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
+Since $L^{-1}=D$ and $J^{-1}=D^{\top}D$, for all $x\in\mathbb{R}^{T}$,
+$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (G.6)
+We next split $b(c)$ into stationary and transient parts. Define the stationary and transient terms as
+$a_{m}^{(P)}:=(P)_{ij}=\frac{1}{n}\quad(m\geq 1),\qquad a_{m}^{(R)}:=(R^{m})_{ij}\quad(m\geq 1),$ (G.7)
+and note $a_{0}=(W^{0})_{ij}=\mathbf{1}\{i=j\}$. Accordingly, write
+$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (G.8)
+where $b^{(0)}$ denotes the (possible) $a_{0}$ contribution,
+$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,c_{k}=\frac{1}{n}\sum_{k=1}^{t-1}c_{k},$ (G.9)
+and
+$b_{t}^{(R)}(c):=\sum_{k=1}^{t-1}(R^{t-k})_{ij}c_{k}=\sum_{m=1}^{t-1}(R^{m})_{ij}\,c_{t-m}.$ (G.10)
+We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
+$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
+$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
+The term $b^{(0)}(c)$ contributes at most an $O(1)$ additive constant, so we ignore it in the asymptotics. By (G.5) and (G.6),
+$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq n\,\|Db^{(P)}(c)\|_{2}^{2}.$
+From (G.9), $(Db^{(P)}(c))_{t}=\frac{1}{n}c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence $\|Db^{(P)}(c)\|_{2}^{2}=(T-1)/n^{2}$ and therefore
+$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{T-1}{n}.$ (G.11)
+By Lemma G.2 given in Section G.2, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
+$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{T-1}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{T}{n}+O(\sqrt{T}).$
+Taking the maximum over $c$ gives
+$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\leq\frac{1}{n}.$ (G.12)
+We derive a lower bound using $c=\mathbf{1}$, i.e., $c$ being all ones. Then
+$(Db)_{t}=b_{t}-b_{t-1}=(W^{t-1})_{ij}=\frac{1}{n}+r_{t-1}.$
+Hence
+$\|Db\|_{2}^{2}=\sum_{t=1}^{T}\left(\frac{1}{n}+r_{t-1}\right)^{2}=\frac{T}{n^{2}}+\frac{2}{n}\sum_{t=1}^{T}r_{t-1}+\sum_{t=1}^{T}r_{t-1}^{2}.$
+Using $|r_{t-1}|\leq\rho^{t-1}$ and summing geometric series, we have
+$\left|\sum_{t=1}^{T}r_{t-1}\right|\leq\sum_{t=1}^{\infty}\rho^{t-1}=\frac{1}{1-\rho}$
+and
+$\sum_{t=1}^{T}r_{t-1}^{2}\leq\sum_{t=1}^{\infty}\rho^{2(t-1)}=\frac{1}{1-\rho^{2}}.$
+Therefore,
+$\|Db\|_{2}^{2}\geq\frac{T}{n^{2}}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}.$ (G.13)
+Next, recall that
+$K_{T}=H_{T}H_{T}^{\top}=\frac{1}{n}J+G_{T},$
+where $J=LL^{\top}$ with $J_{t,s}=\min(t,s)$, and $G_{T}\succeq 0$, and $J^{-1}=D^{\top}D$, where $D$ is the backward difference operator.
+Since $K_{T}$ is positive definite, we use the variational representation
+$b^{\top}K_{T}^{-1}b=\max_{x\in\mathbb{R}^{T}}\left(2b^{\top}x-x^{\top}K_{T}x\right).$
+We lower bound this expression using
+$x_{0}:=nJ^{-1}b=nD^{\top}Db.$
 Then
-$S^{-1}WS\;=\;\begin{bmatrix}1&0\\ 0&T^{\prime}\end{bmatrix}$,
-and taking powers gives
-$W^{t}\;=\;S\begin{bmatrix}1&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\;=\;\underbrace{\mathbf{1}\,\pi^{\top}}_{=\ \Pi}\;+\;\underbrace{S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}}_{=:R^{t}}.$
-To bound $\|R^{t}\|$, fix any matrix norm $\|\cdot\|$.
-Since $\rho(T^{\prime}):=\max\{|\lambda|:\lambda\in\mathrm{spec}(T^{\prime})\}<1$, the Gelfand formula implies that for any $\rho$ with $\rho(T^{\prime})<\rho<1$ there exists $C<\infty$ such that $\|(T^{\prime})^{t}\|\leq C\,\rho^{t}$ for all $t\geq 1$.
+$b^{\top}K_{T}^{-1}b$ $\geq 2b^{\top}x_{0}-x_{0}^{\top}K_{T}x_{0}$
+$=2n\,b^{\top}J^{-1}b-x_{0}^{\top}\left(\tfrac{1}{n}J+G_{T}\right)x_{0}$
+$=n\,b^{\top}J^{-1}b-x_{0}^{\top}G_{T}x_{0}$
+$=n\|Db\|_{2}^{2}-x_{0}^{\top}G_{T}x_{0}.$
+Thus, the required task is to bound the term $x_{0}^{\top}G_{T}x_{0}$. Write
+$W=\frac{1}{n}11^{\top}+R$
+with $1^{\top}R=0$ and $\|R^{t}\|_{2}\leq\rho^{t}$. This gives a decomposition
+$H_{T}=H_{T}^{1}+H_{T}^{R},$
+where $H_{T}^{1}$ is generated by $\frac{1}{n}11^{\top}$ and $H_{T}^{R}$ by $R$. Consequently,
+$G_{T}=H_{T}^{R}(H_{T}^{R})^{\top}.$
+Since $\|R^{t}\|_{2}\leq\rho^{t}$, a geometric-series argument gives (the 2-norm if defined by the first block column of the block lower-triangular matrix $H_{T}^{R}$)
+$\|H_{T}^{R}\|_{2}\leq\frac{1}{1-\rho}.$
+Recall that
+$x_{0}=nD^{\top}Db,\qquad(Db)_{t}=\frac{1}{n}+r_{t-1},\qquad|r_{t}|\leq\rho^{t}.$
+Let $y:=Db$. We then have that
+$(D^{\top}y)_{t}=y_{t}-y_{t+1}\quad(1\leq t\leq T-1),\qquad(D^{\top}y)_{T}=y_{T}.$
+Hence, for $1\leq t\leq T-1$,
+$(x_{0})_{t}=n(y_{t}-y_{t+1})=n(r_{t-1}-r_{t}),$
+and
+$(x_{0})_{T}=ny_{T}=1+nr_{T-1}.$
+Using $|r_{t-1}-r_{t}|\leq|r_{t-1}|+|r_{t}|\leq(1+\rho)\rho^{t-1}$, we obtain
+$\sum_{t=1}^{T-1}(x_{0})_{t}^{2}\leq n^{2}(1+\rho)^{2}\sum_{t=1}^{\infty}\rho^{2(t-1)}=n^{2}\frac{(1+\rho)^{2}}{1-\rho^{2}}=n^{2}\frac{1+\rho}{1-\rho}.$
+Moreover,
+$(x_{0})_{T}^{2}=(1+nr_{T-1})^{2}\leq(1+n)^{2}.$
+Combining these gives
+$\|x_{0}\|_{2}^{2}\leq(1+n)^{2}+n^{2}\frac{1+\rho}{1-\rho},$
+and furthermore
+$x_{0}^{\top}G_{T}x_{0}\leq\frac{(1+n)^{2}}{(1-\rho)^{2}}+n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$ (G.14)
+Combining the bounds of Eq. (G.13) and (G.14) gives
+$b^{\top}K_{T}^{-1}b\geq\frac{T}{n}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}-\frac{(1+n)^{2}}{(1-\rho)^{2}}-n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$
+Since the particular choice of $c=\mathbf{1}$ gives a lower bound, we see that
+$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\geq\frac{1}{n}.$
+This together with (G.12) completes the proof. ∎
+G.2 Bound for the Transient Term
+We next prove a technical lemma required for the proof of Thm. 13. This lemma shows that the transient terms stay bounded independent of $T$.
+Lemma G.2 (Uniform bound for the transient component).
+Let $c\in\{-1,1\}^{T}$ and let $b^{(R)}(c)$ and $K_{T}$ be defined as above in the proof of Thm.
+13.
+There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
+$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
+where $C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}(1-\rho)^{-4}$ and $\rho:=\max_{k\geq 2}|\lambda_{k}|$.
+Proof.
+For the proof we repeatedly use the Schur complement (Gallier, 2010). For $t=1,\dots,T$, let $K_{t}$ be the leading $t\times t$ principal submatrix of $K_{T}$ and write
+$K_{t}=\begin{pmatrix}K_{t-1}&k_{t}\\ k_{t}^{\top}&\kappa_{t}\end{pmatrix},\qquad s_{t}:=\kappa_{t}-k_{t}^{\top}K_{t-1}^{-1}k_{t}>0.$
+For any $x\in\mathbb{R}^{T}$ and $t\in[T]$, define the residual
+$r_{t}(x):=x_{t}-k_{t}^{\top}K_{t-1}^{-1}x_{1:t-1}.$
+A standard block inverse identity with Schur complement gives
+$x^{\top}K_{T}^{-1}x=\sum_{t=1}^{T}\frac{r_{t}(x)^{2}}{s_{t}}.$ (G.15)
+We next lower bound $s_{t}$ uniformly.
+Let $h_{t}^{\top}$ be the $t$th row of $H_{T}$.
+Then $K_{T}$ is the Gram matrix of $\{h_{t}\}_{t=1}^{T}$, and $s_{t}$ equals the squared distance from $h_{t}$ to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$.
+Since $H_{T}$ is block lower-triangular with diagonal block $S=e_{i}^{\top}$, each diagonal component is orthogonal to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$ and we clearly have that
+$s_{t}\geq\|S\|_{2}^{2}=1.$ (G.16)
+Combining (G.15)–(G.16) gives
+$x^{\top}K_{T}^{-1}x\leq\sum_{t=1}^{T}r_{t}(x)^{2}.$ (G.17)
+We now bound the residuals for $x=b^{(R)}(c)$.
+Since $k_{t}^{\top}K_{t-1}^{-1}$ is the minimizer of $\alpha^{\top}\mapsto(x_{t}-\alpha^{\top}x_{1:t-1})^{2}$, for any $\widehat{x}_{t}=\alpha^{\top}x_{1:t-1}$,
+$r_{t}(x)^{2}\leq(x_{t}-\widehat{x}_{t})^{2}.$ (G.18)
+Fix $t\geq 2$ and choose a truncation length
+$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
+Define a predictor $\widehat{b}^{(R)}_{t}$ as
+$\widehat{b}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,b^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii}.$ (G.19)
+We show that the error $e_{t}:=b^{(R)}_{t}-\widehat{b}^{(R)}_{t}$ is $O\!\big(\rho^{m_{t}}/(1-\rho)\big)$. Using (G.10) we may write
+$b^{(R)}_{t}(c)=\sum_{m=1}^{t-1}r_{m}\,c_{t-m},\qquad r_{m}:=(R^{m})_{ij}.$
+Similarly,
+$b^{(R)}_{t-q}(c)=\sum_{m=1}^{t-q-1}r_{m}\,c_{t-q-m}=\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell},$
+and therefore
+$\widehat{b}^{(R)}_{t}=\sum_{q=1}^{m_{t}}\beta_{q}\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell}=\sum_{\ell=2}^{t-1}\left(\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$
+Hence
+$e_{t}=\sum_{\ell=1}^{t-1}\left(r_{\ell}-\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$ (G.20)
+Now use the identity $W^{q}R^{\ell-q}=R^{\ell}$ for all $1\leq q\leq\ell-1$, which follows from $W=P+R$ and $PR=RP=0$. Taking the $(i,j)$ entry,
+$r_{\ell}=(R^{\ell})_{ij}=\sum_{u=1}^{n}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.21)
+Recalling $\beta_{q}=(W^{q})_{ii}$ gives
+$r_{\ell}-\beta_{q}r_{\ell-q}=\sum_{u\neq i}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.22)
+Summing (G.22) over $q=1,\dots,\min\{m_{t},\ell-1\}$ and plugging into (G.20), we see that for all $\ell\geq m_{t}+1$ the coefficient of $c_{t-\ell}$ is a sum of terms $(W^{q})_{iu}(R^{\ell-q})_{uj}$ with $u\neq i$ and $1\leq q\leq m_{t}$.
+Using row-stochasticity ($\sum_{u\neq i}(W^{q})_{iu}\leq 1$) and $|(R^{k})_{uj}|\leq\|R^{k}\|_{2}\leq\rho^{k}$, we obtain
+$\left|r_{\ell}-\sum_{q=1}^{m_{t}}\beta_{q}r_{\ell-q}\right|\leq\sum_{q=1}^{m_{t}}\sum_{u\neq i}(W^{q})_{iu}\,|(R^{\ell-q})_{uj}|\leq\sum_{q=1}^{m_{t}}\rho^{\ell-q}\leq\frac{\rho^{\ell-m_{t}}}{1-\rho}.$
+For $\ell\leq m_{t}$ the bracket in (G.20) is uniformly bounded by
+$|r_{\ell}|+\sum_{q\leq\ell-1}|\beta_{q}||r_{\ell-q}|\leq\rho^{\ell}+\sum_{q\leq\ell-1}\rho^{\ell-q}\leq 1/(1-\rho).$
+Combining these bounds and using $|c_{t-\ell}|=1$ gives
+$|e_{t}|\leq\sum_{\ell=m_{t}+1}^{t-1}\frac{\rho^{\ell-m_{t}}}{1-\rho}\leq\frac{1}{(1-\rho)^{2}}\rho^{m_{t}}\leq\frac{1}{(1-\rho)^{2}}\frac{1}{t}.$ (G.23)
+Finally, by (G.18) and (G.23),
+$r_{t}(b^{(R)}(c))^{2}\leq e_{t}^{2}\leq\frac{1}{(1-\rho)^{4}}\frac{1}{t^{2}}.$ (G.24)
+Using (G.17), summing over $t$ and using the bound (G.24) gives
+$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq\sum_{t=1}^{T}r_{t}(b^{(R)}(c))^{2}\leq\frac{1}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}.$
+This proves the claim. ∎
+Appendix H Extension to Multiple Observed Nodes
+We next give proof of Theorem 15.
+The proof follows exactly the same structure as the proof of Theorem 13: we express the sensitivity as a quadratic form, decompose the dynamics into a stationary component and a transient component, show that the stationary part gives the term linear in $T$, and prove via an auxiliary technical lemma that transient contributions are uniformly bounded in $T$.
+H.1 Proof of Theorem 15
+Proof.
+We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
+$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{mT},$
+and set
+$K_{T}:=H_{T}H_{T}^{\top}.$
+Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
+$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
+hence
+$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (H.1)
+Denote
+$a_{\ell}:=S_{\mathcal{A}}W^{\ell}e_{j}\in\mathbb{R}^{m},\qquad\ell\in\mathbb{N}.$
+Then for $t=1,\dots,T$, the $t$th $m$-block of $b(c)$ satisfies the same convolution identity
+$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}\,c_{k}.$ (H.2)
+We start by decomposing $K_{T}$. By Lemma 9 (Lemma 9 in Appendix C), we can write
+$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (H.3)
+and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|\in(0,1)$.
+Using (H.3) together with the block-Toeplitz structure of $H_{T}$ and the fact that $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$, one obtains the decomposition
+$K_{T}=\frac{1}{n}(J\otimes I_{m})+G_{T},$ (H.4)
+where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$ with $L$ the $T\times T$ lower-triangular all-ones matrix. Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
+$K_{T}\succeq\frac{1}{n}(J\otimes I_{m})$
+which further implies
+$K_{T}^{-1}\preceq n(J^{-1}\otimes I_{m}).$ (H.5)
+Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
+$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
+so that $J^{-1}=D^{\top}D$ and for all $x\in\mathbb{R}^{T}$,
+$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (H.6)
+We next split $b(c)$ into stationary and transient parts, exactly as in the proof of Thm. 13. Define the stationary and transient terms as
+$a_{\ell}^{(P)}:=S_{\mathcal{A}}Pe_{j}=\frac{1}{n}\mathbf{1}_{m}\quad(\ell\geq 1),\qquad a_{\ell}^{(R)}:=S_{\mathcal{A}}R^{\ell}e_{j}\quad(\ell\geq 1),$ (H.7)
+and $a_{0}=S_{\mathcal{A}}e_{j}$. Accordingly, write
+$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (H.8)
+for $t=1,\dots,T$,
+$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,\mathbf{1}_{m}\,c_{k}=\frac{1}{n}\mathbf{1}_{m}\sum_{k=1}^{t-1}c_{k},$ (H.9)
+and
+$b_{t}^{(R)}(c):=\sum_{m=1}^{t-1}S_{\mathcal{A}}R^{m}e_{j}\,c_{t-m}.$ (H.10)
+We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
+$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
+$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
+By (H.5) and (H.6),
+$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}$ $\leq n\,(b^{(P)}(c))^{\top}(J^{-1}\otimes I_{m})\,b^{(P)}(c)=n\,\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}.$
+From (H.9), we have $(Db^{(P)}(c))_{t}=\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence
+$\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}=\sum_{t=2}^{T}\left\|\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}\right\|_{2}^{2}=\sum_{t=2}^{T}\frac{m}{n^{2}}=\frac{m(T-1)}{n^{2}}.$
 Therefore
-$\|R^{t}\|\;=\;\left\|S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\right\|\;\leq\;\|S\|\,\|S^{-1}\|\,\|(T^{\prime})^{t}\|\;\leq\;\big(\|S\|\,\|S^{-1}\|\;C\big)\,\rho^{t}.$
-Setting $M:=\|S\|\,\|S^{-1}\|\,C$ gives the claim. ∎
-Appendix E Proof Theorem 13
+$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{m(T-1)}{n}.$ (H.11)
+By Lemma H.1, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
+$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{m(T-1)}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{m}{n}T+O(\sqrt{T}).$
+The constants given in the theorem statement are obtained by combining the $O(1)$ term with the $\sqrt{\frac{m(T-1)}{n}}$-term. ∎
+H.2 Bound for the Transient Term
+We next prove the required technical lemma which shows that the transient terms stay bounded independent of $T$ also for the multi-node observation.
+Lemma H.1.
+Let $c\in\{-1,1\}^{T}$ and let $\widetilde{b}^{(R)}(c)$ and $\widetilde{K}_{T}$ be defined as above in the proof of Thm.
+15.
+There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
+$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq C_{\mathrm{mix}}(W),$ (H.12)
+where one may take
+$C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}\,m\,(1-\rho)^{-4},\qquad\rho:=\max_{k\geq 2}|\lambda_{k}(W)|.$
 Proof.
-First, from the compact SVD we see that for any matrix $A$, the product $A^{+}A$ is the orthogonal projector onto $\mathrm{row}(A)$.
-Set $P_{T}:=H_{T}^{+}H_{T}$ to be the orthogonal projection onto row space of $H_{T}$.
-Let $c\in\{\pm 1\}^{T}$ and define $z:=c\otimes e_{j}\in\mathbb{R}^{nT}$.
-Our strategy is to study the sensitivity
-$\Delta_{T}(c,j\to i)=\|P_{T}z\|_{2}$
-and derive a bound for it that is independent of $c$.
-Because $W$ is primitive and row-stochastic, by Lemma 11 there exist constants $M>0$ and $\rho\in(0,1)$ such that
-$W^{t}=\Pi+R^{t},\qquad\Pi:=\mathbf{1}\pi^{\top},\qquad\|R^{t}\|_{2}\leq M\rho^{t}\quad\forall t\geq 0.$ (E.1)
-The $(t,s)$ block of $H_{T}$ (with $0\leq s\leq t\leq T-1$) is given by
-$H_{T}(t,s)=e_{i}^{\top}W^{t-s}=e_{i}^{\top}\Pi+e_{i}^{\top}R^{t-s}=\pi^{\top}+e_{i}^{\top}R^{t-s}.$
-Define the stationary part of $H_{T}$ as $\bar{H}_{T}$ such that blocks of $\bar{H}_{T}$ are given as
-$\bar{H}_{T}(t,s)=\pi^{\top}\quad\text{ for }s\leq t,\qquad\bar{H}_{T}(t,s)=0\quad\text{for }s>t.$
-Furthermore, define the transient part
-$K_{T}:=H_{T}-\bar{H}_{T}.$
+The proof follows the same Schur-complement argument as Lemma G.2, with the only difference that now the vectors live in $\mathbb{R}^{mT}$.
+For $t=1,\dots,T$, let $\widetilde{K}_{t}$ be the leading $(mt)\times(mt)$ principal submatrix of $\widetilde{K}_{T}$ and write it in block form
+$\widetilde{K}_{t}=\begin{pmatrix}\widetilde{K}_{t-1}&\widetilde{k}_{t}\\ \widetilde{k}_{t}^{\top}&\widetilde{\kappa}_{t}\end{pmatrix},\qquad\widetilde{s}_{t}:=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}\succ 0,$
+where $\widetilde{\kappa}_{t}\in\mathbb{R}^{m\times m}$ and $\widetilde{s}_{t}\in\mathbb{R}^{m\times m}$. For any $x\in\mathbb{R}^{mT}$ define the residual vector
+$\widetilde{r}_{t}(x):=x_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}x_{1:t-1}\in\mathbb{R}^{m},$
+where $x_{t}\in\mathbb{R}^{m}$ denotes the $t$th block of $x$. A standard block inverse identity with Schur complement gives
+$x^{\top}\widetilde{K}_{T}^{-1}x=\sum_{t=1}^{T}\widetilde{r}_{t}(x)^{\top}\widetilde{s}_{t}^{-1}\widetilde{r}_{t}(x).$ (H.13)
+We next lower bound $\widetilde{s}_{t}$.
+Let $\widetilde{h}_{t}^{\top}$ be the $t$th block-row of $\widetilde{H}_{T}$.
+Then, $\widetilde{K}_{T}$ is the Gram matrix of $\{\widetilde{h}_{t}\}_{t=1}^{T}$.
+Here $\widetilde{s}_{t}$ admits a geometric interpretation analogous to the scalar case.
+Since $\widetilde{K}_{T}$ is the Gram matrix of the block-rows $\{\widetilde{h}_{1},\dots,\widetilde{h}_{T}\}$ of $\widetilde{H}_{T}$, the Schur complement
+$\widetilde{s}_{t}=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}$
+is the Gram matrix of the component of $\widetilde{h}_{t}$ orthogonal to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$.
+Equivalently, for any $z\in\mathbb{R}^{m}$, $z^{\top}\widetilde{s}_{t}z$ equals the squared distance of the scalar row $z^{\top}\widetilde{h}_{t}$ from the span of $\{z^{\top}\widetilde{h}_{1},\dots,z^{\top}\widetilde{h}_{t-1}\}$, and $\widetilde{s}_{t}$ equals the squared-distance operator (in $\mathbb{R}^{m}$) from $\widetilde{h}_{t}$ to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$.
+Since $\widetilde{H}_{T}$ is block lower-triangular with diagonal block $US_{\mathcal{A}}$, and $US_{\mathcal{A}}$ has orthonormal rows (because $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$ and $U$ is orthogonal), the diagonal block is orthogonal to the past block-rows, implying
+$\widetilde{s}_{t}\succeq I_{m}\quad\Rightarrow\quad\widetilde{s}_{t}^{-1}\preceq I_{m}.$ (H.14)
+Combining (H.13)–(H.14) gives
+$x^{\top}\widetilde{K}_{T}^{-1}x\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(x)\|_{2}^{2}.$ (H.15)
+We now bound the residuals for $x=\widetilde{b}^{(R)}(c)$.
+As in the proof of Lemma G.2, the residual $\widetilde{r}_{t}(x)$ is the error of the best linear predictor of $x_{t}$ from the past $x_{1:t-1}$.
+Hence for any (possibly suboptimal) predictor $\widehat{x}_{t}$ measurable w.r.t.
+the past,
+$\|\widetilde{r}_{t}(x)\|_{2}^{2}\leq\|x_{t}-\widehat{x}_{t}\|_{2}^{2}.$ (H.16)
+Fix $t\geq 2$ and set the truncation length
+$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
+We use the same predictor as in the proof of Lemma G.2, applied component-wise:
+$\widehat{\widetilde{b}}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,\widetilde{b}^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii},$ (H.17)
+for an arbitrary but fixed $i\in\mathcal{A}$ (any $i$ works; only $|\beta_{q}|\leq 1$ is used).
+Using same techniques as in the proof of Lemma G.2, i.e., using (i) $W^{q}R^{\ell-q}=R^{\ell}$, (ii) row-stochasticity, and (iii) $\|R^{k}\|_{2}\leq\rho^{k}$, gives the uniform tail bound
+$\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}\leq\frac{\sqrt{m}}{(1-\rho)^{2}}\,\frac{1}{t}.$ (H.18)
+Here the extra factor $\sqrt{m}$ comes from bounding an $\ell_{2}$-norm over $m$ coordinates by $\sqrt{m}$ times a coordinate-wise bound.
+Combining (H.16) and (H.18), we obtain
+$\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\,\frac{1}{t^{2}}.$ (H.19)
+Using (H.15), summing over $t$ and using the bound (H.19) gives
+$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}\,m.$
+This proves the claim. ∎
+Appendix I Extension to the Case When Observers’ Noise Does not Contribute (SecLDP)
+We next extend the proof technique of Theorem 15 to the SecLDP setting, where the adversary observing a set $\mathcal{A}$ can remove the Gaussian noise injected by nodes in $\mathcal{A}$.
+Equivalently, one deletes the corresponding columns in each time-block column of the system matrix.
+I.1 Proof of Theorem 16
+We simply sketch the proof as it has exactly the same structure as the proof of Theorem 15.
+In the SecLDP setting, the adversary observing $\mathcal{A}$ can remove the Gaussian noises injected by nodes in $\mathcal{A}$.
+This is modeled by right-multiplying each time-block column of the system matrix by
+$P_{\mathcal{A}^{c}}:=I_{n}-\sum_{i\in\mathcal{A}}e_{i}e_{i}^{\top},\qquad\bar{H}_{T}:=H_{T}(I_{T}\otimes P_{\mathcal{A}^{c}}),\qquad\bar{K}_{T}:=\bar{H}_{T}\bar{H}_{T}^{\top}.$
+For $j\notin\mathcal{A}$, $(I_{T}\otimes P_{\mathcal{A}^{c}})(c\otimes e_{j})=c\otimes e_{j}$, hence: $\bar{b}(c):=\bar{H}_{T}(c\otimes e_{j})=H_{T}(c\otimes e_{j})=b(c)$.
+Therefore the SecLDP sensitivity is
+$\bar{\Delta}_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}\bar{K}_{T}^{+}\,b(c),$
+and the difference to Theorem 15 is in the modified Gram matrix $\bar{K}_{T}$.
+We decompose $W=P+R$ as in the single- and multi-node proofs, with $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$, $PR=RP=0$, and $\|R^{t}\|_{2}\leq\rho^{t}$.
+This gives a decomposition $\bar{H}_{T}=\bar{H}_{T}^{(P)}+\bar{H}_{T}^{(R)}$ and, expanding,
+$\bar{K}_{T}=\bar{K}_{T}^{(PP)}+\bar{K}_{T}^{(RR)}+\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)},\qquad\bar{K}_{T}^{(PP)}:=\bar{H}_{T}^{(P)}(\bar{H}_{T}^{(P)})^{\top},\ \ \bar{K}_{T}^{(RR)}:=\bar{H}_{T}^{(R)}(\bar{H}_{T}^{(R)})^{\top}.$
+Unlike in the proof of Theorem 15, the cross terms $\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)}$ do not vanish in general because of the right-multiplication by $P_{\mathcal{A}^{c}}$ (we cannot use the orthogonality $\mathbf{1}^{\top}R=0$).
+We keep these terms explicitly and control them by Cauchy–Schwarz: for any $x$ and any $\eta>0$,
+$\big|x^{\top}(\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)})x\big|\leq\eta\,x^{\top}\bar{K}_{T}^{(PP)}x+\eta^{-1}x^{\top}\bar{K}_{T}^{(RR)}x,$
+so the cross terms can be absorbed into the stationary and transient terms.
+The stationary term is computed explicitly. Using $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $P_{\mathcal{A}^{c}}\mathbf{1}=\mathbf{1}_{\mathcal{A}^{c}}$, one obtains the Toeplitz form
+$\bar{K}_{T}^{(PP)}=\frac{n-m}{n^{2}}\,(J\otimes\mathbf{1}_{m}\mathbf{1}_{m}^{\top}),\qquad J_{t,s}=\min(t,s),$
+i.e., compared to the proof of Theorem 15, we get the coefficient (nm)/n(n-m)/n because only the $\mathcal{A}^{c}$ noise coordinates remain.
+Consequently, the same backward-difference argument ($J^{-1}=D^{\top}D$) used in (H.11) gives the SecLDP linear term with slope $\frac{m}{n-m}$ (instead of $\frac{m}{n}$).
+Finally, all terms involving $R$ (the pure transient contribution and the cross-term contribution) are controlled by the exactly the same Schur-complement technique as Lemma G.2 and Lemma H.1, giving a uniform bound independent of $T$:
+$(b^{(R)}(c))^{\top}\bar{K}_{T}^{+}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
+and similarly for the cross terms. Putting the stationary and transient terms together gives the claim.
+Appendix J Proofs for Section 5 (Utility Analysis)
+J.1 From Absolute Spectral Gap to the Consensus Parameter $p$
+We prove the standard contraction inequality and its implication for the parameter $p$ that is needed in the unified decentralized SGD analysis of (Koloskova et al., 2020).
+Lemma J.1.
+Let $W\in\mathbb{R}^{n\times n}$ be symmetric and doubly-stochastic.
+Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be its eigenvalues.
+For any $X\in\mathbb{R}^{d\times n}$, define $\bar{X}:=X\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $\rho=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}$.
+Then
+$\|XW-\bar{X}\|_{F}^{2}\;\leq\;\rho^{2}\,\|X-\bar{X}\|_{F}^{2}.$ (J.1)
+Consequently, the assumption
+$\|XW-\bar{X}\|_{F}^{2}\leq(1-p)\|X-\bar{X}\|_{F}^{2}$ (J.2)
+needed in the analysis of (Koloskova et al., 2020) holds with $p=1-\rho^{2}$.
+Proof.
+Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$.
+Then $\bar{X}=XP$ and, since $W\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}W=\mathbf{1}^{\top}$, we have $WP=PW=P$.
+Hence
+$XW-\bar{X}=XW-XP=X(W-P)=(X-\bar{X})(W-P),$
+because $\bar{X}(W-P)=XP(W-P)=X(PW-P)=0$.
+Now, WPW-P is symmetric with eigenvalues 0 (along $\mathbf{1}$) and $\lambda_{2}(W),\ldots,\lambda_{n}(W)$ on $\mathbf{1}^{\perp}$, so $\|W-P\|_{2}=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}=\rho$ (since $W$ is symmetric).
 Thus
-$H_{T}=\bar{H}_{T}+K_{T}.$
-We clearly have that
-$\mathrm{row}(\bar{H}_{T})=\{(\alpha_{0}\pi,\alpha_{1}\pi,\dots,\alpha_{T-1}\pi)\in\mathbb{R}^{nT}:\alpha_{t}\in\mathbb{R}\}$
-and we can take as an orthonormal basis for $\mathrm{row}(\bar{H}_{T})$ for example $\{e_{1}\otimes\pi/\|\pi\|_{2},\ldots,e_{T}\otimes\pi/\|\pi\|_{2}\}.$ Therefore the orthogonal projection onto $\mathrm{row}(\bar{H}_{T})$ is given by
-$\bar{P}_{T}z=\sum_{\ell=1}^{T}\langle e_{\ell}\otimes\pi/\|\pi\|_{2},c\otimes e_{j}\rangle=c\;\otimes\;\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi.$
-and we have that
-$\|\bar{P}_{T}z\|_{2}^{2}=\sum_{t=0}^{T-1}\left\|c_{t}\,\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi\right\|_{2}^{2}=T\,\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
-We next study the transient part and show using standard perturbation theory for projection matrices that the distance between the projectors $P_{T}$ and $\bar{P}_{T}$ satisfies
-$\|P_{T}-\bar{P}_{T}\|_{2}=O(1/\sqrt{T}).$
-This rate is derived using the Wedin theorem which states that the distance between the row spaces $\mathrm{row}(H_{T})$ and $\mathrm{row}(\bar{H}_{T})$ is a function of $\|K_{T}\|_{2}$ (since the perturbation $H_{T}-\bar{H}_{T}=K_{T}$) and the smallest non-zero singular value of $\bar{H}_{T}$, $\sigma_{\min}(\bar{H}_{T})$.
-From the Toeplitz structure it follows that clearly $\|K_{T}\|_{2}$ equals the 2-norm of the first column block of $K_{T}$.
-Moreover, from the transient behavior stated in Eq.
-(E.1) and a geometric sum it follows that $\|K_{T}\|_{2}=O(1)$.
-We see that
-$\bar{H}_{T}=U\otimes\pi^{\top}$
-for
-$U=\begin{pmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{pmatrix}$.
-The singular values of a Kronecker product $A\otimes B$ are the products of the singular values of $A$ and $B$. Thus, the non-zero singular values of $\bar{H}_{T}$ are:
-$\sigma_{k}(\bar{H}_{T})=\sigma_{m}(U)\cdot\sigma_{l}(\pi^{\top}).$
-Since $\pi^{\top}$ has rank $1$, $\sigma_{1}(\pi^{\top})=\|\pi^{\top}\|_{2}=\|\pi\|_{2}.$ We also know that the smallest singular value of $U$ is given by
-$\sigma_{\min}(U)=\sqrt{T}.$
-Thus, $\sigma_{\min}(\bar{H}_{T})=\sqrt{T}\|\pi\|_{2}$, and by the Wedin theorem we have that
-$\|P_{T}-\bar{P}_{T}\|_{2}\leq\frac{2\|K_{T}\|_{2}}{\sigma_{\min}(\bar{H}_{T})}=O(1/\sqrt{T}).$
-Let $\mathcal{E}_{T}=P_{T}-\bar{P}_{T}$, so $\|\mathcal{E}_{T}\|_{2}=O(1/\sqrt{T})$.
-We expand $\|P_{T}z\|_{2}^{2}$ using $P_{T}=\bar{P}_{T}+\mathcal{E}_{T}$. Recall $z=c\otimes e_{j}$ and $\|z\|_{2}=\sqrt{T}$:
-$\|P_{T}z\|_{2}^{2}=\|(\bar{P}_{T}+\mathcal{E}_{T})z\|_{2}^{2}=\|\bar{P}_{T}z\|_{2}^{2}+2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)+\|\mathcal{E}_{T}z\|_{2}^{2}.$ (E.2)
-Using a norm bound on $\mathcal{E}_{T}$, we have
-$|2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)|$ $\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}z\|_{2}$
-$\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}\|_{2}\|z\|_{2}$
-$\leq 2\left(\sqrt{T}\frac{\pi_{j}}{\|\pi\|_{2}}\right)\cdot O(1/\sqrt{T})\cdot\sqrt{T}=O(\sqrt{T}).$
-Furthermore,
-$\|\mathcal{E}_{T}z\|_{2}^{2}\leq\|\mathcal{E}_{T}\|_{2}^{2}\|z\|_{2}^{2}=O(1/T)\cdot T=O(1).$
-Substituting these bounds back into Eq. (E.2):
-$\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
-This holds for any choice of $c\in\{\pm 1\}^{T}$. The maximum is thus:
-$(\Delta_{T}(j\to i))^{2}=\max_{c\in\{\pm 1\}^{T}}\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
-Dividing by $T$ and taking the limit:
-$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\lim_{T\to\infty}\left(\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(1/\sqrt{T})\right)=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
+$\|XW-\bar{X}\|_{F}=\|(X-\bar{X})(W-P)\|_{F}\leq\|X-\bar{X}\|_{F}\,\|W-P\|_{2}=\rho\,\|X-\bar{X}\|_{F}.$
+Using the relation $1-\rho^{2}=(1+\rho)(1-\rho)$ we directly get the following auxiliary result.
+Lemma J.2.
+Let $W$ be symmetric and doubly-stochastic and define the absolute spectral gap
+$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$
+With $p=1-\rho^{2}$,
+$\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right),\;\;\frac{1}{p^{2}}=O\!\left(\frac{1}{\gamma^{2}}\right).$ (J.3)
+J.2 Deriving the Utility Bound from Results of (Koloskova et al., 2020)
+We derive Theorem 17 by plugging our privacy calibration into a strongly-convex convergence bound by (Koloskova et al., 2020).
+More specifically, we use the following template bound (a specialization of the unified analysis to $\tau=1$ and strongly convex objectives).
+Lemma J.3 (Koloskova et al. 2020).
+Assume each $f_{i}$ is $L$-smooth and $\mu$-strongly convex.
+Let $\tau=1$ (full-batch gradients).
+Assume unbiased stochastic gradients with variance proxy $\bar{\sigma}^{2}$ and heterogeneity $\bar{\zeta}^{2}$.
+Assume the consensus condition (J.2) holds with parameter $p\in(0,1]$.
+Then, for a constant step size $\eta\simeq 1/L$, the averaged iterate $\bar{\theta}^{t}$ satisfies
+$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{\theta}^{t})-f^{\star}\right]\;\leq\;\widetilde{O}\!\left(\frac{\bar{\sigma}^{2}}{n\mu T}\;+\;\frac{L(\bar{\zeta}^{2}+p\bar{\sigma}^{2})(1-p)}{\mu^{2}p^{2}T^{2}}\;+\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{p}\exp\!\left(-c_{0}\,\frac{\mu p}{L}T\right)\right),$ (J.4)
+[…]
+J.3 Proof of Theorem 17
+Proof.
+To specialize Lemma J.3 to our setting, we start by identifying the gradient noise as the DP noise.
+Namely, under (5.1) with full gradients and $\xi_{i}^{t}\sim\mathcal{N}(0,\sigma_{\mathrm{DP}}^{2}I_{d})$, the only stochasticity is the DP noise, and hence
+$\sigma_{i}^{2}:=\mathbb{E}\|\xi_{i}^{t}\|_{2}^{2}=d\,\sigma_{\mathrm{DP}}^{2},\qquad\bar{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}=d\,\sigma_{\mathrm{DP}}^{2}.$
+Next, we adjust Gaussian noise to $(\varepsilon,\delta)$-DP guarantees by (see, e.g., Dwork and Roth, 2014)
+$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\sqrt{2\log(1.25/\delta)}}{\varepsilon}.$
+As shown in the main text, using Thm. 15, we have
+$(\Delta_{T}(j\to i))^{2}\;\leq\;\frac{T}{n}+\frac{2}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{2}{\gamma^{4}}$ (J.5)
+and by adjusting the Gaussian noise via the standard condition (Dwork and Roth, 2014)
+$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\,\sqrt{2\log(1.25/\delta)}}{\varepsilon}$ (J.6)
+we get
+$\sigma_{\mathrm{DP}}^{2}\;\lesssim\;\frac{\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right).$ (J.7)
+For symmetric doubly-stochastic $W$, Lemma J.1 gives $p=1-\gamma^{2}$ and Lemma J.2 implies $p\geq\gamma$ and $1/p=O(1/\gamma)$, $1/p^{2}=O(1/\gamma^{2})$.
+We now bound each term in (J.4).
+(i) Leading term. Using (J.7),
+$\frac{\bar{\sigma}^{2}}{n\mu T}$ $=\widetilde{O}\!\left(\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n^{2}\,\mu}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{2}\sqrt{T}}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{4}T}\right)$ (J.8)
+(ii) Heterogeneity term. We split:
+$\frac{L(\bar{\zeta}^{2}+p\bar{\sigma}^{2})(1-p)}{\mu^{2}p^{2}T^{2}}\;\leq\;\frac{L\bar{\zeta}^{2}}{\mu^{2}p^{2}T^{2}}\;+\;\frac{L\bar{\sigma}^{2}}{\mu^{2}pT^{2}},$
+since $(1-p)\leq 1$ and $(p\bar{\sigma}^{2})(1-p)/p^{2}\leq\bar{\sigma}^{2}/p$. Using $p\geq\gamma$, we have that
+$\frac{L\bar{\zeta}^{2}}{\mu^{2}p^{2}T^{2}}$ $\leq\frac{L\bar{\zeta}^{2}}{\mu^{2}\gamma^{2}T^{2}}$ (J.9)
+and
+$\frac{L\bar{\sigma}^{2}}{\mu^{2}pT^{2}}$ $=\widetilde{O}\!\left(\frac{L}{\mu^{2}\gamma T^{2}}\cdot\frac{d\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right)\right)$
+$=\widetilde{O}\!\left(\frac{dL\log(1/\delta)}{\varepsilon^{2}\mu^{2}}\cdot\frac{1}{n\gamma T}\right).$ (J.10)
+(iii) Transient term. Using $p\geq\gamma$,
+$\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{p}\exp\!\left(-c_{0}\frac{\mu p}{L}T\right)\;\leq\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{\gamma}\exp\!\left(-c\,\gamma\frac{\mu}{L}T\right)$ (J.11)
+for some constant $c>0$.
+Combining (J.8), (J.9) and (J.11) gives the statement of Theorem 17. ∎